### v1

# An Adaptive Online Smoother with Closed-Form Solutions and Information-Theoretic Lag Selection for Conditional Gaussian Nonlinear Systems

Abstract
Data assimilation (DA) combines partial observations with a dynamical model to improve state estimation. Filter-based DA uses only past and present data and is the prerequisite for real-time forecasts. Smoother-based DA exploits both past and future observations. It aims to fill in missing data, provide more accurate estimations, and develop high-quality datasets. However, the standard smoothing procedure requires using all historical state estimations, which is storage-demanding, especially for high-dimensional systems. This paper develops an adaptive-lag online smoother for a large class of complex dynamical systems with strong nonlinear and non-Gaussian features, which has important applications to many real-world problems. The adaptive lag allows the DA to utilize only observations within a nearby window, significantly reducing computational storage. Online lag adjustment is essential for tackling turbulent systems, where temporal autocorrelation varies significantly over time due to intermittency, extreme events, and nonlinearity. Based on the uncertainty reduction in the estimated state, an information criterion is developed to systematically determine the adaptive lag. Notably, the mathematical structure of these systems facilitates the use of closed analytic formulae to calculate the online smoother and the adaptive lag, avoiding empirical tunings as in ensemble-based DA methods. The adaptive online smoother is applied to studying three important scientific problems. First, it helps detect online causal relationships between state variables. Second, its advantage of computational storage is illustrated via Lagrangian DA, a high-dimensional nonlinear problem. Finally, the adaptive smoother advances online parameter estimation with partial observations, emphasizing the role of the observed extreme events in accelerating convergence.
1 Introduction
Complex turbulent nonlinear dynamical systems (CTNDSs) have broad applications across various fields [1, 2, 3, 4]. These systems are characterized by their high dimensionality and multiscale structures, with strong nonlinear interactions occurring between state variables at different spatiotemporal scales. Extreme events, intermittency, and non-Gaussian probability density functions (PDFs) are commonly observed in these systems [5, 6, 7]. State estimation is essential for parameter estimation, prediction, optimal control, and generating complete datasets [8, 9, 10]. However, the turbulent nature of dynamics can amplify small errors in model structure, spatiotemporal solutions, or initial conditions when relying solely on forecasts. Data assimilation (DA), which integrates observations with system dynamics, is widely used to improve state estimation [11, 12, 13, 14, 15]. Given the inevitable uncertainty in state estimation, especially for the unobserved variables in a CTNDS, probabilistic state approaches via Bayesian inference are natural choices. The model provides a prior distribution, while observations inform the likelihood. They are combined to form the posterior distribution for state estimation.
DA can be classified into two categories based on when observational data is incorporated. Filtering uses observations only up to the current time. Serving as the initialization, filter-based state estimation is the prerequisite for real-time forecasts. In contrast, smoothing [15, 16, 17, 18, 19] leverages data from the entire observation period, including future data, which makes it highly effective for optimal state estimation in offline data postprocessing. This helps to fill missing values, minimize bias, and create complete datasets [20]. With the extra information from future observations, smoothing often produces more accurate and less uncertain state estimates than filtering. When the system dynamics and observational mappings are linear, with additive Gaussian noise, the corresponding filtering and smoothing methods are the Kalman filter and the Rauch-Tung-Striebel (RTS) smoother, respectively [16, 21, 22], where the posterior distribution can be computed using closed-form analytical solutions.
Due to the intrinsic nonlinear dynamics and non-Gaussian statistics of CTNDSs, analytic solutions for DA are rarely available. As a result, various numerical and approximate methods have been developed, including the ensemble Kalman filter/smoother and the particle (or sequential Monte Carlo) filter/smoother [14, 15, 23, 24, 25, 26, 27]. These methods are widely used but often face tremendous computational costs, especially in high-dimensional systems [28], which limits the number of particles or ensemble members, potentially causing biases and numerical instabilities [29, 30, 31]. Empirical tuning techniques, such as noise inflation, localization, and resampling, are widely used in practice to mitigate these issues [23, 32, 33, 34]. However, these ad hoc tuning methods are usually quite challenging to implement systematically. Closed-form analytic solutions for DA are thus highly desirable, as they improve computational efficiency, stability, and accuracy, especially in capturing non-Gaussian features, including intermittency and extreme events. They also facilitate theoretical analysis of error and uncertainty in state estimation.
Instead of refining DA schemes directly, computational challenges in state estimation can be addressed by developing approximate models that yield analytic solutions for the posterior distribution. While linear approximations allow for standard methods like the Kalman filter or RTS smoother, linearizing a strongly nonlinear system often leads to biases and instabilities. An alternative is a recently developed class of nonlinear systems that includes many turbulent models in geophysics, fluids, engineering, and neuroscience [35, 36, 37, 38, 39]. Despite their nonlinear dynamics and non-Gaussian statistics, the conditional distributions of unobserved state variables given observations are Gaussian (which is precisely the posterior distribution in the DA context), leading to the term conditional Gaussian nonlinear systems (CGNSs). The CGNS framework allows the use of closed analytic formulae for solving these conditional distributions, helping develop efficient algorithms for filtering, smoothing, and sampling without the ad hoc tuning often needed in ensemble-based DA methods. It also facilitates rigorous analysis of these methods. Additionally, CGNSs have been utilized as surrogate models in various applications, including DA, prediction, preconditioning, and machine learning [39, 40, 41, 42].
The standard smoother-based state estimation procedure involves executing a forward pass for filtering across the entire observational period, followed by a backward pass for smoothing [16, 17]. However, the standard offline smoother requires storing the filter solution for the entire duration before initiating the backward pass, which requires substantial computational storage, particularly in high-dimensional systems. Due to the wide application of smoother-based state estimation, it is of practical importance to develop a computationally efficient and accurate algorithm that significantly reduces storage.
This paper presents a forward-in-time online smoother algorithm with adaptive lags for the CGNS framework, eliminating the need for a full backward pass. The online smoother sequentially updates the current state as new observations become available. By doing so, it effectively addresses the computational storage issue. While online schemes exist for the RTS smoother and ensemble-based methods, the CGNS online smoother has several unique advantages. First, despite the intrinsic nonlinearity of the underlying dynamics, closed analytic formulae are available to compute the nonlinear online smoother. These analytic formulae provide precise and accurate solutions, which avoid numerical and sampling errors as in ensemble-based methods. Second, due to the turbulent nature of the system, observations influence the estimated state only within a short time window, which enhances computational efficiency and reduces storage needs. Third, different from fixed-lag smoothers [17, 27, 43, 44], the lag in the CGNS smoother is adaptively determined. Online lag adjustment is essential for studying turbulent systems, where temporal autocorrelation varies significantly over time due to intermittency, extreme events, and nonlinearity. A fixed lag usually either overuses storage (if the lag is overestimated) or introduces a large bias (if the lag is underestimated). In contrast, an adaptive lag optimizes the use of data and computational storage. Finally, the adaptive lag is systematically determined using an information criterion based on the uncertainty reduction in the posterior distribution [38, 45]. It emphasizes the importance of the posterior uncertainty and differs from some of the existing adaptive lag selection criteria that rely solely on the posterior mean [46]. As closed analytic formulae are available for posterior distributions, the information gain can be computed efficiently and accurately.
The adaptive online smoother for the CGNS framework is applied to studying three important scientific problems. First, the online update of the smoother estimate allows for quantification of the improvement in state estimation by incorporating future information. It facilitates revealing causal dependence between state variables. A nonlinear dyad model with strong non-Gaussian features is utilized for such a study. Second, the CGNS framework is applied to Lagrangian data assimilation (DA), which is a high-dimensional nonlinear problem that has a significant storage requirement [42, 47, 48, 49]. The online smoother allows for the estimation of the unobserved flow states based on Lagrangian tracers. The study highlights the role of the online smoother in reducing computational storage. Finally, the online smoother facilitates developing an online parameter estimation algorithm with partial observations. It helps reveal the role of the observed intermittent extreme events in advancing parameter estimation.
The remainder of this paper is organized as follows. Section 2 introduces the CGNS modeling framework, including the equations for the offline optimal nonlinear filter and smoother state estimation. Section 3 presents the adaptive online smoother. In Section 4, the application of the adaptive online smoother to the three key problems is demonstrated. Section 5 includes the conclusion. The appendices contain detailed analysis and proofs.
2 The Conditional Gaussian Nonlinear System Modelling Framework
Throughout this paper, boldface letters are exclusively used to denote vectors for the sake of mathematical clarity. In this regard, we use lowercase boldface letters to denote column vectors, while Uppercase boldface letters denote matrices. The only exception to this rule is $\mathbf{W}$ (with some subscript or superscript), which denotes a Wiener process. Although in this work, this always corresponds to a column vector, we instead use an uppercase letter due to literary tradition.
Let $t$ denote the time variable, with $t\in[0,T]$, where $T>0$ is allowed to be infinite. Let $(\Omega,\mathcal{F},\mathbb{P})$ be a complete probability space and $\{\mathcal{F}_{t}\}_{t\in[0,T]}$ to be a filtration of sub-$\sigma$-algebras of $(\Omega,\mathcal{F})$, which we assume is augmented (i.e., complete and right-continuous). For every filtration there exists a smallest such augmented filtration refining $\{\mathcal{F}_{t}\}_{t\in[0,T]}$ (known as its completion), so this is without loss of generality. We let $(\mathbf{x}(t,\omega),\mathbf{y}(t,\omega))$, for $t\in[0,T]$ and $\omega\in\Omega$ (for the rest of this work we drop the event or sample space dependence for notational simplicity, but it is always implied), be a partially observable $(S,\mathcal{A})$-valued stochastic process, where $\mathbf{x}$ is the observable component, while $\mathbf{y}$ is the unobservable component. The theory that follows can be applied mutatis mutandis to any partially observable stochastic process that takes values over a measurable space $(S,\mathcal{A})$ where $S$ is a separable Hilbert space (finite-dimensional or not) over a complete scalar ground field and $\mathcal{A}$ is a $\sigma$-algebra of $S$, but for this work it suffices to consider complex-valued finite-dimensional processes with respect to the usual Euclidean inner product. As such, we let $S=\mathbb{C}^{k+l}$ and $\mathcal{A}=\mathcal{B}_{\mathbb{C}^{k+l}}\equiv\mathcal{B}_{\mathbb{R}^{2(k+l)}}$, with $\mathbf{x}$ being a $k$-dimensional vector and $\mathbf{y}$ an $l$-dimensional one, where $\mathcal{B}_{\mathbb{R}^{2(k+l)}}$ is the Borel $\sigma$-algebra of $\mathbb{R}^{2(k+l)}$, since $\mathrm{dim}_{\mathbb{R}}(\mathbb{C}^{k+l})=2(k+l)$. We assume that
$(\mathbf{x},\mathbf{y})=\big{(}(x_{1}(t),\ldots,x_{k}(t),y_{1}(t),\ldots,y_{l}(t)),\mathcal{F}_{t}\big{)},\ t\in[0,T],$
meaning the partially observable random process is (jointly) adapted to the filtration $\{\mathcal{F}_{t}\}_{t\in[0,T]}$, i.e., for all times $t\in[0,T]$ the random vector defined by $(\mathbf{x}(t,\boldsymbol{\cdot})^{\mathtt{T}},\mathbf{y}(t,\boldsymbol{\cdot})^{\mathtt{T}})^{\mathtt{T}}:\Omega\to S$ is an $(\mathcal{F}_{t};\mathcal{A})$-measurable function. Specifically, this implies that the natural filtration of $\mathcal{F}$ with respect to $\{\mathbf{x}(s)\}_{s\leq t}$, which is the sub-$\sigma$-algebra generated by the observable processes for times $s\leq t$ and is defined by
$\mathcal{F}_{t}^{\mathbf{x}}:=\sigma(\{\mathbf{x}(s)\}_{s\leq t})=\{\mathbf{x}(s)^{-1}[A]=\mathbf{x}(s,\boldsymbol{\cdot})^{-1}[A]:A\in\mathcal{A},\ s\leq t\},$
satisfies $\mathcal{F}_{t}^{\mathbf{x}}\subseteq\mathcal{F}_{t}$, since $\mathbf{x}$ is adapted to the filtration $\{\mathcal{F}_{t}\}_{t\in[0,T]}$ by construction and by definition $\mathcal{F}_{t}^{\mathbf{x}}$ is the smallest such filtration. We call this natural filtration the observable $\sigma$-algebra (at time $t$) for the remainder of this work.
It is known that, at each time instant $t\in[0,T]$, the optimal estimate in the mean-square sense of some $\mathcal{F}_{t}$-measurable function $h(\mathbf{x}(t),\mathbf{y}(t))$ of $\mathbf{x}$ and $\mathbf{y}$, on the basis of the observations up to time $t$, $\{\mathbf{x}(s)\}_{s\leq t}$, is exactly its conditional expectation conditioned on the observable $\sigma$-algebra at time $t$, $\mathbb{E}\left[\mathbf{h}(\mathbf{x}(t),\mathbf{y}(t))|\mathcal{F}_{t}^{\mathbf{x}}\right]$. This is known as the a-posteriori mean and this assertion of optimality rests on the tacit assumption that $\mathbb{E}\left[\left\|\mathbf{h}(\mathbf{x}(t),\mathbf{y}(t))\right\|_{2}^{2}\right]$ is finite, where $\left\|\boldsymbol{\cdot}\right\|_{2}$ denotes the usual Euclidean norm over $\mathbb{C}^{\mathrm{dim}(\mathbf{h})}$ [35, 36]. Usually $\mathbf{h}$ is a function of the unobserved process $\mathbf{y}$, and in this work we exclusively use $\mathbf{h}=\mathbf{y}$ as to recover the optimal filter and smoother conditional Gaussian statistics of the hidden process when conditioning on the observations.
The goal of optimal state estimation is to characterize the posterior states using a system of stochastic differential equations (SDEs), known as the optimal nonlinear filter or smoother equations, depending on whether we condition on current observations or the entire observational period, respectively. In general, without making specific assumptions about the structure of the processes $\mathbf{h}$ and $\mathbf{x}$, determining $\mathbb{E}\left[\mathbf{h}|\mathcal{F}_{t}^{\mathbf{x}}\right]$ is challenging. Nevertheless, the conditional Gaussian nonlinear system (CGNS) has a key advantage in its conditional Gaussianity of the posterior distributions, which can be written down using closed analytic formulae. Note that the classical Kalman-Bucy filter [21] is the simplest example of a CGNS [36]. Despite the conditional Gaussianity, these coupled systems remain highly nonlinear and the associated marginal and joint distributions are highly non-Gaussian, which allows the systems to capture many realistic turbulent features.
2.1 Conditionally Gaussian Processes
In its most general form, a conditional Gaussian system of processes consists of two diffusion-type processes defined by the following system of stochastic differentials given in Itô form [36, 38, 50]:
$\mathrm{d}\mathbf{x}(t)$ $=\left[\boldsymbol{\Lambda}^{\mathbf{x}}(t,\mathbf{x})\mathbf{y}(t)+\mathbf{f}^{\mathbf{x}}(t,\mathbf{x})\right]\mathrm{d}t+\boldsymbol{\Sigma}_{1}^{\mathbf{x}}(t,\mathbf{x})\mathrm{d}\mathbf{W}_{1}(t)+\boldsymbol{\Sigma}_{2}^{\mathbf{x}}(t,\mathbf{x})\mathrm{d}\mathbf{W}_{2}(t),$ (2.1)
$\mathrm{d}\mathbf{y}(t)$ $=\left[\boldsymbol{\Lambda}^{\mathbf{y}}(t,\mathbf{x})\mathbf{y}(t)+\mathbf{f}^{\mathbf{y}}(t,\mathbf{x})\right]\mathrm{d}t+\boldsymbol{\Sigma}_{1}^{\mathbf{y}}(t,\mathbf{x})\mathrm{d}\mathbf{W}_{1}(t)+\boldsymbol{\Sigma}_{2}^{\mathbf{y}}(t,\mathbf{x})\mathrm{d}\mathbf{W}_{2}(t),$ (2.2)
where
$\mathbf{W}_{1}=\big{(}(W_{11}(t),\ldots,W_{1d}(t)),\mathcal{F}_{t}\big{)}\quad\text{and}\quad\mathbf{W}_{2}=\big{(}(W_{21}(t),\ldots,W_{2r}(t)),\mathcal{F}_{t}\big{)},$
are two mutually independent complex-valued Wiener processes (i.e., both their real and imaginary parts are mutually independent real-valued Wiener processes) and almost every path of $\mathbf{x}$ and $\mathbf{y}$ is in $C^{0}([0,T];\mathbb{C}^{k})$ and $C^{0}([0,T];\mathbb{C}^{l})$, respectively. The elements of the vector- and matrix-valued functions of multiplicative factors ($\boldsymbol{\Lambda}^{\mathbf{x}},\boldsymbol{\Lambda}^{\mathbf{y}}$), forcings ($\mathbf{f}^{\mathbf{x}},\mathbf{f}^{\mathbf{y}}$), and noise feedbacks ($\boldsymbol{\Sigma}_{1}^{\mathbf{x}},\boldsymbol{\Sigma}_{2}^{\mathbf{x}},\boldsymbol{\Sigma}_{1}^{\mathbf{y}},\boldsymbol{\Sigma}_{2}^{\mathbf{y}}$) are assumed to be nonanticipative (adapted) functionals over the measurable time-function cylinder
$(C^{0,k}_{T},\mathscr{B}^{k}_{T}):=\big{(}[0,T]\times C^{0}([0,T];\mathbb{C}^{k}),\mathcal{B}([0,T])\otimes\mathcal{B}(C^{0}([0,T];\mathbb{C}^{k}))\big{)},$
where $\otimes$ denotes the tensor-product $\sigma$-algebra on the underlying product space, i.e.,
$\mathscr{B}^{k}_{T}=\mathcal{B}([0,T])\otimes\mathcal{B}(C^{0}([0,T];\mathbb{C}^{k}))=\sigma\left(\big{\{}A\times B:A\in\mathcal{B}([0,T]),B\in\mathcal{B}(C^{0}([0,T];\mathbb{C}^{k}))\big{\}}\right),$
with $\mathcal{B}(C^{0}([0,T];\mathbb{C}^{k}))$ being the $\sigma$-algebra generated by the topology of compact convergence on the space of continuous functions from $[0,T]$ to $\mathbb{C}^{k}$, $C^{0}([0,T];\mathbb{C}^{k})$. This means that if $g(t,\mathbf{x})$ denotes a standard element of these functionals, for each $t\in[0,T]$, it is $(\mathcal{F}_{t}^{\mathbf{x}};\mathbb{C})$-measurable. It is important to emphasize here the fact that in a CGNS, the unobservable component $\mathbf{y}$ enters the dynamics in a conditionally linear manner, whereas the observable process $\mathbf{x}$ can enter into the coefficients of both equations in any $\mathcal{F}^{\mathbf{x}}_{t}$-measurable way.
Many CTNDSs fit into the CGNS modelling framework. Some well-known classes of these systems are physics-constrained nonlinear stochastic models (for example the noisy versions of the Lorenz models, low-order models of Charney-DeVore flows, and a paradigm model for topographic mean flow interaction) [51, 52], stochastically coupled reaction-diffusion models used in neuroscience and ecology (for example the stochastically coupled FitzHugh-Nagumo models and the stochastically coupled SIR epidemic models), and spatiotemporally multiscale models for turbulence, fluids, and geophysical flows (for example the Boussinesq equations with noise and the stochastically forced rotating shallow-water equation) [38]. This modelling framework has also been exploited to develop realistic low-order stochastic models for the Madden-Julian oscillation (MJO) and Arctic sea ice [41, 53].
In addition to modeling many natural phenomena, the CGNS framework and its closed analytic DA formulae have been applied to study many theoretical and practical problems. It has been utilized to develop a nonlinear LDA algorithm, allowing rigorous analysis to study model error and uncertainty, as well as recovery of turbulent ocean flows with noisy observations from Lagrangian tracers [42, 47, 48, 54]. The analytically solvable DA scheme has also been applied to the prediction and state estimation of the non-Gaussian intermittent time series of the MJO and the monsoon intraseasonal variabilities, in addition to the filtering of the stochastic skeleton model for the MJO [37, 41, 55, 56, 57, 58]. Notably, the efficient DA procedure also helps develop a rapid algorithm to solve the high-dimensional Fokker-Planck equation [59, 60]. Worth highlighting is that the ideas of the CGNS modelling framework and the associated DA procedures have also been applied to develop cheap exactly-solvable forecast models in dynamic stochastic superresolution of sparsely observed turbulent systems [61, 62], build stochastic superparameterization for geophysical turbulence [63, 64, 65, 66], and design efficient multiscale DA schemes via blended particle filters for high-dimensional chaotic systems [67].
A set of sufficient regularity conditions needs to be assumed a-priori so that the main result of the CGNS framework can be established. We enforce the same set of assumptions as in the work of Andreou & Chen [68], which we outline in Appendix A, with the rationale behind the adoption of each one being provided there (and in the references therein). These conditions are then sufficient enough to show that the posterior distribution of the unobserved variables, when conditioning on the observational data, is Gaussian, as stated in the following theorem; this is exactly why the system in (2.1)–(2.2) is called a CGNS of equations. For the following, we abuse notation for clarity and instead write $\big{(}\boldsymbol{\cdot}\big{|}\mathbf{x}(s),s\leq t\big{)}$ to indicate the fact that we are conditioning on the observable $\sigma$-algebra at time $t$, $\big{(}\boldsymbol{\cdot}\big{|}\mathcal{F}_{t}^{\mathbf{x}}\big{)}$.
Theorem 2.1 (Conditional Gaussianity).
Let $(\mathbf{x}(t),\mathbf{y}(t))$ satisfy (2.1)–(2.2) and assume that the regularity conditions (1)(6) found in Appendix A hold. Additionally, assume that the initial conditional distribution $\mathbb{P}\big{(}\mathbf{y}(0)\leq\boldsymbol{\alpha}_{0}\big{|}\mathbf{x}(0)\big{)}$ [^1] The event $\{\mathbf{y}(s)\leq\boldsymbol{\alpha}=(\alpha_{1},\ldots,\alpha_{l})^{\mathtt{T}}\}$ is to be understood coordinate-wise, i.e., $\{y_{1}(s)\leq\alpha_{1},\ldots,y_{l}(s)\leq\alpha_{l}\}$, with the linear partial ordering assumed on $\mathbb{C}$, i.e., $y_{j}(s)\leq\alpha_{j}\Leftrightarrow\big{(}\mathrm{Re}(y_{j}(s))<\mathrm{Re}(\alpha_{j})\big{)}\vee\big{(}\mathrm{Re}(y_{j}(s))=\mathrm{Re}(\alpha_{j})\wedge\mathrm{Im}(y_{j}(s))\leq\mathrm{Im}(\alpha_{j})\big{)}$. be ($\mathbb{P}$-a.s.) Gaussian, $\mathcal{N}_{l}(\boldsymbol{\mu}_{\text{f}}(0),\mathbf{R}_{\text{f}}(0))$, and mutually independent from the Wiener processes $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$, where
$\boldsymbol{\mu}_{\text{f}}(0):=\mathbb{E}\left[\mathbf{y}(0)\big{|}\mathbf{x}(0)\right]\quad\text{and}\quad\mathbf{R}_{\text{f}}(0):=\mathbb{E}\left[(\mathbf{y}(0)-\boldsymbol{\mu}_{\text{f}}(0))(\mathbf{y}(0)-\boldsymbol{\mu}_{\text{f}}(0))^{\dagger}\big{|}\mathbf{x}(0)\right],$
with $\boldsymbol{\cdot}^{\dagger}$ denoting the Hermitian transpose operator. Furthermore, assume $\mathbb{P}\left(\mathrm{tr}(\mathbf{R}_{\text{f}}(0))<+\infty\right)=1$, where $\mathrm{tr}(\boldsymbol{\cdot})$ denotes the trace operator, meaning that the initial estimation mean-square error between $\mathbf{y}(0)$ and $\boldsymbol{\mu}_{\text{f}}(0)$ is almost surely finite. Then, for any $t_{j}$ such that $0\leq t_{1}<t_{2}<\cdots<t_{n}\leq t$, with $t\in[0,T]$, and $\boldsymbol{\alpha}_{1},\ldots,\boldsymbol{\alpha}_{n}\in\mathbb{C}^{l}$, the conditional distribution
$\mathbb{P}\big{(}\mathbf{y}(t_{1})\leq\boldsymbol{\alpha}_{1},\ldots,\mathbf{y}(t_{n})\leq\boldsymbol{\alpha}_{n}\big{|}\mathbf{x}(s),s\leq t\big{)},$
is ($\mathbb{P}$-a.s.) Gaussian.
Proof.
This is Theorem 12.6 in Liptser & Shiryaev [36], which is the multi-dimensional analog of Theorem 11.1. For the analogous result in the case of discrete time, see Theorem 13.3, with the respective sufficient assumptions given in Subchapter 13.2.1. Thorough details are also provided in Kolodziej [69] for the continuous-time case. The proof, regardless of continuous- or discrete-time, uses the conditional characteristic function method and a conditional version of the law-uniqueness theorem [69, 70]. ∎
2.2 Analytically Solvable Filter and Smoother Posterior Distributions
With Theorem 2.1 established, it is then possible to yield the optimal nonlinear filter state estimation equations as showed in the following theorem, where the subscript “ f ” is used to denote the filter conditional Gaussian statistics, which appropriately stands for filter. The filter conditional Gaussian statistics are also known as the filter posterior mean and filter posterior covariance under the Bayesian inference dynamics framework.
Theorem 2.2 (Optimal Nonlinear Filter State Estimation Equations).
Let the assumptions of Theorem 2.1 and the additional regularity conditions (7)(9), which are outlined in Appendix A, to hold. Then for any $t\in[0,T]$ the $\mathcal{F}^{\mathbf{x}}_{t}$-measurable Gaussian statistics of the Gaussian conditional distribution
$\mathbb{P}\big{(}\mathbf{y}(t)\leq\boldsymbol{\alpha}\big{|}\mathbf{x}(s),s\leq t\big{)}\overset{{\,\rm d}}{\sim}\mathcal{N}_{l}(\boldsymbol{\mu}_{\text{{f}}}(t),\mathbf{R}_{\text{{f}}}(t)),$
where
$\boldsymbol{\mu}_{\text{{f}}}(t):=\mathbb{E}\left[\mathbf{y}(t)\big{|}\mathbf{x}(s),s\leq t\right]\quad\text{and}\quad\mathbf{R}_{\text{{f}}}(t):=\mathbb{E}\left[(\mathbf{y}(t)-\boldsymbol{\mu}_{\text{{f}}}(t))(\mathbf{y}(t)-\boldsymbol{\mu}_{\text{{f}}}(t))^{\dagger}\big{|}\mathbf{x}(s),s\leq t\right],$
are the unique continuous solutions of the system of optimal nonlinear filter equations:
${\,\rm d}\boldsymbol{\mu}_{\text{{f}}}(t)$ $=(\boldsymbol{\Lambda}^{\mathbf{y}}\boldsymbol{\mu}_{\text{{f}}}+\mathbf{f}^{\mathbf{y}}){\,\rm d}t+(\boldsymbol{\Sigma}^{\mathbf{y}}\circ\boldsymbol{\Sigma}^{\mathbf{x}}+\mathbf{R}_{\text{{f}}}(\boldsymbol{\Lambda}^{\mathbf{x}})^{\dagger})(\boldsymbol{\Sigma}^{\mathbf{x}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})^{-1}({\,\rm d}\mathbf{x}(t)-(\boldsymbol{\Lambda}^{\mathbf{x}}\boldsymbol{\mu}_{\text{{f}}}+\mathbf{f}^{\mathbf{x}}){\,\rm d}t),$ (2.3)
${\,\rm d}\mathbf{R}_{\text{{f}}}(t)$ $=\big{(}\boldsymbol{\Lambda}^{\mathbf{y}}\mathbf{R}_{\text{{f}}}+\mathbf{R}_{\text{{f}}}(\boldsymbol{\Lambda}^{\mathbf{y}})^{\dagger}+\boldsymbol{\Sigma}^{\mathbf{y}}\circ\boldsymbol{\Sigma}^{\mathbf{y}}-(\boldsymbol{\Sigma}^{\mathbf{y}}\circ\boldsymbol{\Sigma}^{\mathbf{x}}+\mathbf{R}_{\text{{f}}}(\boldsymbol{\Lambda}^{\mathbf{x}})^{\dagger})(\boldsymbol{\Sigma}^{\mathbf{x}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})^{-1}(\boldsymbol{\Sigma}^{\mathbf{x}}\circ\boldsymbol{\Sigma}^{\mathbf{y}}+\boldsymbol{\Lambda}^{\mathbf{x}}\mathbf{R}_{\text{{f}}})\big{)}{\,\rm d}t,$ (2.4)
with initial conditions $\boldsymbol{\mu}_{\text{{f}}}(0)=\mathbb{E}\left[\mathbf{y}(0)\big{|}\mathbf{x}(0)\right]$ and $\mathbf{R}_{\text{{f}}}(0)=\mathbb{E}\left[(\mathbf{y}(0)-\boldsymbol{\mu}_{\text{{f}}}(0))(\mathbf{y}(0)-\boldsymbol{\mu}_{\text{{f}}}(0))^{\dagger}\big{|}\mathbf{x}(0)\right]$, where the noise interactions through the Gramians (with respect to rows) are defined as
$(\boldsymbol{\Sigma}^{\mathbf{x}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})(t,\mathbf{x}):=\boldsymbol{\Sigma}_{1}^{\mathbf{x}}(t,\mathbf{x})\boldsymbol{\Sigma}_{1}^{\mathbf{x}}(t,\mathbf{x})^{\dagger}+\boldsymbol{\Sigma}_{2}^{\mathbf{x}}(t,\mathbf{x})\boldsymbol{\Sigma}_{2}^{\mathbf{x}}(t,\mathbf{x})^{\dagger},$
$(\boldsymbol{\Sigma}^{\mathbf{y}}\circ\boldsymbol{\Sigma}^{\mathbf{y}})(t,\mathbf{x}):=\boldsymbol{\Sigma}_{1}^{\mathbf{y}}(t,\mathbf{x})\boldsymbol{\Sigma}_{1}^{\mathbf{y}}(t,\mathbf{x})^{\dagger}+\boldsymbol{\Sigma}_{2}^{\mathbf{y}}(t,\mathbf{x})\boldsymbol{\Sigma}_{2}^{\mathbf{y}}(t,\mathbf{x})^{\dagger},$
$(\boldsymbol{\Sigma}^{\mathbf{x}}\circ\boldsymbol{\Sigma}^{\mathbf{y}})(t,\mathbf{x}):=\boldsymbol{\Sigma}_{1}^{\mathbf{x}}(t,\mathbf{x})\boldsymbol{\Sigma}_{1}^{\mathbf{y}}(t,\mathbf{x})^{\dagger}+\boldsymbol{\Sigma}_{2}^{\mathbf{x}}(t,\mathbf{x})\boldsymbol{\Sigma}_{2}^{\mathbf{y}}(t,\mathbf{x})^{\dagger},\quad(\boldsymbol{\Sigma}^{\mathbf{y}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})(t,\mathbf{x}):=(\boldsymbol{\Sigma}^{\mathbf{x}}\circ\boldsymbol{\Sigma}^{\mathbf{y}})(t,\mathbf{x})^{\dagger}.$
Furthermore, if the initial covariance matrix $\mathbf{R}_{\text{{f}}}(0)$ is positive-definite ($\mathbb{P}$-a.s.), then all the matrices $\mathbf{R}_{\text{{f}}}(t)$, for $t\in[0,T]$, remain positive-definite ($\mathbb{P}$-a.s.).
Proof.
This is Theorem 12.7 in Liptser & Shiryaev [36], which is the multi-dimensional analog of Theorems 12.1 and 12.3. For the analogous result in the case of discrete time, see Theorem 13.4, which outlines the corresponding optimal recursive nonlinear filter difference equations, with the respective sufficient assumptions given in Subchapter 13.2.1. Thorough details are also provided in Kolodziej [69] for the continuous-time case. For a martingale-free proof see Andreou & Chen [68] (although it requires the additional assumptions (10)(11) found in Appendix A, as to pass from the discrete-time filter to the continuous one via a formal limit). ∎
The form of the filter mean equation can be intuitively explained within the Bayesian inference framework for DA. The first two terms on the right-hand side of (2.3), namely $\boldsymbol{\Lambda}^{\mathbf{y}}\boldsymbol{\mu}_{\text{f}}+\mathbf{f}^{\mathbf{y}}$, represent the forecast mean, derived from the process of the unobserved variables in (2.2). The remaining terms account for correcting the posterior mean state, incorporating information from partial observations. The matrix factor in front of the innovation, or pre-fit measurement residual ${\,\rm d}\mathbf{x}(t)-(\boldsymbol{\Lambda}^{\mathbf{x}}\boldsymbol{\mu}_{\text{{f}}}+\mathbf{f}^{\mathbf{x}}){\,\rm d}t$, is analogous to the Kalman gain in classical Kalman filter theory: ${(\boldsymbol{\Sigma}^{\mathbf{y}}\circ\boldsymbol{\Sigma}^{\mathbf{x}}+\mathbf{R}_{\text{{f}}}(\boldsymbol{\Lambda}^{\mathbf{x}})^{\dagger})(\boldsymbol{\Sigma}^{\mathbf{x}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})^{-1}}$. This factor determines the weight of the observations when updating the model-predicted state. Even in cases where the variables $\mathbf{y}$ of the unobserved process do not explicitly depend on $\mathbf{x}$, such as in the Lagrangian DA, the observational process (2.1) still couples the observed and unobserved components. This coupling allows the observations to impact state estimation and correct the forecast. Finally, observe that the equation driving the evolution of the covariance tensor $\mathbf{R}_{\text{f}}$ is a random Riccati equation [71, 72, 73], since the coefficients depend on the observable random variables $\mathbf{x}$.
As already stated, the CGNS framework also enjoys closed analytic formulae for the recovery of the optimal smoother state estimation. The smoother posterior distribution at time $t\in[0,T]$ exploits the observational information in the entire period $[0,T]$ and therefore it allows for a more accurate, less uncertain, and unbiased estimated state compared to the filter solution. Solving the optimal nonlinear smoother equations requires applying a forward pass (filtering) from $t=0$ to $t=T$, which is then followed by a backward pass (smoothing) from $t=T$ to $t=0$ [15, 16, 17, 18, 19]. In the theorem that follows, the optimal nonlinear smoother equations, which run backwards in time, are showed, with the subscript “ s ” is denoting the smoother conditional Gaussian statistics, which appropriately stands for smoother. The smoother conditional Gaussian statistics are also known as the smoother posterior mean and smoother posterior covariance under the Bayesian inference dynamics framework.
Theorem 2.3 (Optimal Nonlinear Smoother State Estimation Backward Equations).
Let the assumptions of Theorem 2.1 and the additional regularity conditions (7)(9), which are outlined in Appendix A, to hold. In addition, assume
$\mathbb{P}\left(\underset{t\in[0,T]}{\mathrm{inf}}\{\mathrm{det}(\mathbf{R}_{\text{{f}}}(t))\}>0\right)=1,$ (2.5)
and define the following auxiliary matrices:
$\mathbf{A}(t,\mathbf{x})$ $:=\boldsymbol{\Lambda}^{\mathbf{y}}(t,\mathbf{x})-(\boldsymbol{\Sigma}^{\mathbf{y}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})(t,\mathbf{x})(\boldsymbol{\Sigma}^{\mathbf{x}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})^{-1}(t,\mathbf{x})\boldsymbol{\Lambda}^{\mathbf{x}}(t,\mathbf{x})\in\mathbb{C}^{l\times l},$ (2.6)
$\mathbf{B}(t,\mathbf{x})$ $:=(\boldsymbol{\Sigma}^{\mathbf{y}}\circ\boldsymbol{\Sigma}^{\mathbf{y}})(t,\mathbf{x})-(\boldsymbol{\Sigma}^{\mathbf{y}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})(t,\mathbf{x})(\boldsymbol{\Sigma}^{\mathbf{x}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})^{-1}(t,\mathbf{x})(\boldsymbol{\Sigma}^{\mathbf{x}}\circ\boldsymbol{\Sigma}^{\mathbf{y}})(t,\mathbf{x})\in\mathbb{C}^{l\times l}.$ (2.7)
Then the system of optimal nonlinear smoother equations is given by the following backward differential equations:
$\overleftarrow{{\,\rm d}\boldsymbol{\mu}_{\text{{s}}}}$ $=-\left(\boldsymbol{\Lambda}^{\mathbf{y}}\boldsymbol{\mu}_{\text{{s}}}+\mathbf{f}^{\mathbf{y}}-\mathbf{B}\mathbf{R}_{\text{{f}}}^{-1}(\boldsymbol{\mu}_{\text{{f}}}-\boldsymbol{\mu}_{\text{{s}}})\right){\,\rm d}t+(\boldsymbol{\Sigma}^{\mathbf{y}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})(\boldsymbol{\Sigma}^{\mathbf{x}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})^{-1}\left(\overleftarrow{{\,\rm d}\mathbf{x}}+(\boldsymbol{\Lambda}^{\mathbf{x}}\boldsymbol{\mu}_{\text{{s}}}+\mathbf{f}^{\mathbf{x}}){\,\rm d}t\right),$ (2.8)
$\overleftarrow{{\,\rm d}\mathbf{R}_{\text{{s}}}}$ $=-\left((\mathbf{A}+\mathbf{B}\mathbf{R}_{\text{{f}}}^{-1})\mathbf{R}_{\text{{s}}}+\mathbf{R}_{\text{{s}}}(\mathbf{A}+\mathbf{B}\mathbf{R}_{\text{{f}}}^{-1})^{\dagger}-\mathbf{B}\right){\,\rm d}t,$ (2.9)
for $T\geq t\geq 0$, with $t$ running backward. The backward-arrow notation on the left-hand side of (2.8)–(2.9) and in $\overleftarrow{{\,\rm d}\mathbf{x}}$ on right-hand side of (2.8) is to be understood as
$\overleftarrow{{\,\rm d}\boldsymbol{\mu}_{\text{{s}}}}:=\lim_{\Delta t\to 0}\left(\boldsymbol{\mu}_{\text{{s}}}(t)-\boldsymbol{\mu}_{\text{{s}}}(t+\Delta t)\right),\quad\overleftarrow{{\,\rm d}\mathbf{R}_{\text{{s}}}}:=\lim_{\Delta t\to 0}\left(\mathbf{R}_{\text{{s}}}(t)-\mathbf{R}_{\text{{s}}}(t+\Delta t)\right),\quad\overleftarrow{{\,\rm d}\mathbf{x}}:=\lim_{\Delta t\to 0}\left(\mathbf{x}(t)-\mathbf{x}(t+\Delta t)\right),$
for $\boldsymbol{\mu}_{\text{{s}}}(t)$ and $\mathbf{R}_{\text{{s}}}(t)$ being the conditional mean and covariance of the smoother posterior distribution at time $t$, which is Gaussian,
$\mathbb{P}\big{(}\mathbf{y}(t)\big{|}\mathbf{x}(s),s\in[0,T]\big{)}\overset{{\,\rm d}}{\sim}\mathcal{N}_{l}(\boldsymbol{\mu}_{\text{{s}}}(t),\mathbf{R}_{\text{{s}}}(t)).$
In other words, the notation $\overleftarrow{\frac{{\,\rm d}\boldsymbol{\cdot}}{{\,\rm d}t}}$ corresponds to the negative of the usual derivative, which means the system in (2.8)–(2.9) is meant to be solved backward over $[0,T]$. The “starting” values for the smoother posterior statistics, $\left(\boldsymbol{\mu}_{\text{{s}}}(T),\mathbf{R}_{\text{{s}}}(T)\right)$, are the same as those of the corresponding filter estimates at the endpoint $t=T$, $\left(\boldsymbol{\mu}_{\text{{f}}}(T),\mathbf{R}_{\text{{f}}}(T)\right)$.
Proof.
This is Theorem 12.10 in Liptser & Shiryaev [36]. For the analogous result in the case of discrete time, see Theorem 13.12, which outlines the corresponding optimal recursive nonlinear smoother backward difference equations, with the respective sufficient assumptions given in Subchapters 13.2.1 and 13.3.8. For a martingale-free proof see Andreou & Chen [68] (although it requires the additional assumptions (10)(12) found in Appendix A, as to pass from the discrete-time smoother to the continuous one via a formal limit). ∎
3 Optimal Online Smoother
The standard smoothing procedure involves a forward pass using a filtering method followed by a backward pass to obtain the optimal smoother state estimation [15, 16, 17, 18, 19]. While this forward-backward algorithm is effective for offline data postprocessing, in many practical situations, new observations become available sequentially. Recalculating the smoother’s posterior Gaussian statistics from scratch each time is computationally and storage-intensive. Therefore, an online version of the optimal nonlinear smoother, requiring only forward-in-time model updates, is highly desirable, as it can update estimated states as new data arrive. While particle filters and ensemble methods offer approximate state estimates for general CTNDSs, the CGNS framework provides exact, closed-form analytic solutions for deriving an online smoother. The online optimal nonlinear smoother not only efficiently updates the estimates but also facilitates the understanding of the mathematical and numerical properties of the system.
3.1 Optimal Online Forward-In-Time Discrete Smoother
We begin by adopting a time discretization scheme, determined by the rate at which observational data arrive, similar to numerical integration. Although our framework allows for irregular observation intervals, for simplicity, we assume the data arrive at a regular, uniform rate, which is common in practice. Thus, observations are available sequentially at times $t_{0}=0<t_{1}<t_{2}<\cdots<t_{n}<\cdots<+\infty$, where $\Delta t_{j}\equiv\Delta t=t_{j+1}-t_{j},$ for all $j$. We also assume that $\Delta t$ is sufficiently small to ensure stability and consistency of the discrete-time schemes for numerical integration of the CGNS equations, as well as of the optimal nonlinear filter and smoother equations, thus ensuring their convergence [68].
In what follows, the superscript notation $\boldsymbol{\cdot}^{\,j}$ is used to denote the discrete approximation to the continuous form of the respective vector or matrix functional when evaluated on $t_{j}$, for example $\boldsymbol{\Lambda}^{\mathbf{x},j}:=\boldsymbol{\Lambda}^{\mathbf{x}}(t_{j},\mathbf{x}(t_{j}))$. We write the CGNS of equations (2.1)–(2.2) in a discrete fashion using the Euler-Maruyama scheme [74, 75, 76]. Then the time discretization of (2.1)–(2.2) is simply given by
$\mathbf{x}^{j+1}$ $=\mathbf{x}^{j}+\left(\boldsymbol{\Lambda}^{\mathbf{x},j}\mathbf{y}^{j}+\mathbf{f}^{\mathbf{x},j}\right)\Delta t+\boldsymbol{\Sigma}_{1}^{\mathbf{x},j}\sqrt{\Delta t}\boldsymbol{\epsilon}_{1}^{j}+\boldsymbol{\Sigma}_{2}^{\mathbf{x},j}\sqrt{\Delta t}\boldsymbol{\epsilon}_{2}^{j},$ (3.1)
$\mathbf{y}^{j+1}$ $=\mathbf{y}^{j}+\left(\boldsymbol{\Lambda}^{\mathbf{y},j}\mathbf{y}^{j}+\mathbf{f}^{\mathbf{y},j}\right)\Delta t+\boldsymbol{\Sigma}_{1}^{\mathbf{y},j}\sqrt{\Delta t}\boldsymbol{\epsilon}_{1}^{j}+\boldsymbol{\Sigma}_{2}^{\mathbf{y},j}\sqrt{\Delta t}\boldsymbol{\epsilon}_{2}^{j},$ (3.2)
where $\boldsymbol{\epsilon}_{1}^{j}$ and $\boldsymbol{\epsilon}_{2}^{j}$ are mutually independent complex standard Gaussian random noises. The explicit nature and order of the temporal discretization used here is sufficient for the CGNS framework we are working with, meaning implicit or higher-order discretization schemes are not needed [68].
Given a series of realizations for the observable $\mathbf{x}$, $\big{\{}\mathbf{x}^{0},\mathbf{x}^{1},\ldots,\mathbf{x}^{n}\big{\}}$, where $\mathbf{x}^{j}$ was obtained $\Delta t$ time units after $\mathbf{x}^{j-1}$ for $j=1,\ldots,n$, we let $\boldsymbol{\mu}_{\text{s}}^{j,n}$ and $\mathbf{R}_{\text{s}}^{j,n}$ denote the discrete smoother posterior mean and the discrete smoother posterior covariance, respectively, when evaluated at time $t=t_{j}$ and conditioned on this realization of $\mathbf{x}$ up to time $t=t_{n}$, where $0\leq j\leq n$ (see Theorem 2.3). In other words, we define
$\boldsymbol{\mu}_{\text{{s}}}^{j,n}:=\mathbb{E}\left[\mathbf{y}^{j}\big{|}\mathbf{x}^{s},s\leq n\right],\quad\mathbf{R}_{\text{{s}}}^{j,n}:=\mathrm{Cov}\left(\mathbf{y}^{j},\mathbf{y}^{j}\big{|}\mathbf{x}^{s},s\leq n\right)=\mathbb{E}\left[(\mathbf{y}^{j}-\boldsymbol{\mu}_{\text{{s}}}^{j,n})(\mathbf{y}^{j}-\boldsymbol{\mu}_{\text{{s}}}^{j,n})^{\dagger}\big{|}\mathbf{x}^{s},s\leq n\right].$
Notice how we explicitly note the dependence of the smoother state estimates on the length of the observational period. As is known by the discrete-time counterpart of Theorem 2.1 (see Theorem 13.3 in Liptser & Shiryaev [36]), the smoother posterior distribution $\mathbb{P}\big{(}\mathbf{y}^{j}\big{|}\mathbf{x}^{s},s=0,\ldots,n\big{)}$ is ($\mathbb{P}$-a.s.) conditionally Gaussian, $\mathcal{N}_{l}(\boldsymbol{\mu}_{\text{{s}}}^{j,n},\mathbf{R}_{\text{{s}}}^{j,n})$, and it is possible to show that the conditional mean $\boldsymbol{\mu}_{\text{{s}}}^{j,n}$ and conditional covariance $\mathbf{R}_{\text{{s}}}^{j,n}$ of the discrete smoother at time step $t_{j}$ when conditioning up to the $n$-th observation satisfy the following recursive backward difference equations for $n\in\mathbb{N}$ and $j=0,1,\ldots,n-1$ [68] (under regularity assumptions (1)(12) in Appendix A and all other assumed conditions):
$\boldsymbol{\mu}_{\text{{s}}}^{j,n}$ $=\boldsymbol{\mu}_{\text{{f}}}^{j}+\mathbf{E}^{j}\left(\boldsymbol{\mu}_{\text{{s}}}^{j+1,n}-(\mathbf{I}_{l\times l}+\boldsymbol{\Lambda}^{\mathbf{y},j}\Delta t)\boldsymbol{\mu}_{\text{{f}}}^{j}-\mathbf{f}^{\mathbf{y},j}\Delta t\right)+\mathbf{F}^{j}\left(\mathbf{x}^{j+1}-\mathbf{x}^{j}-(\boldsymbol{\Lambda}^{\mathbf{x},j}\boldsymbol{\mu}_{\text{{f}}}^{j}+\mathbf{f}^{\mathbf{x},j})\Delta t\right),$ (3.3)
$\mathbf{R}_{\text{{s}}}^{j,n}$ $=\mathbf{R}_{\text{f}}^{j}+\mathbf{E}^{j}\left(\mathbf{R}_{\text{{s}}}^{j+1,n}(\mathbf{E}^{j})^{\dagger}-(\mathbf{I}_{l\times l}+\boldsymbol{\Lambda}^{\mathbf{y},j}\Delta t)\mathbf{R}_{\text{f}}^{j}\right)-\mathbf{F}^{j}\boldsymbol{\Lambda}^{\mathbf{x},j}\mathbf{R}_{\text{f}}^{j}\Delta t,$ (3.4)
where $\mathbf{I}_{l\times l}$ is the $l\times l$ identity matrix and the auxiliary matrices appearing in (3.3)–(3.4) are given up to leading-order $O(\Delta t)$, since we are using the Euler-Maruyama scheme to temporally discretize the CGNS of equations, as
$\mathbf{E}^{j}$ $:=\mathbf{I}_{l\times l}+\left[(\boldsymbol{\Sigma}^{\mathbf{y}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})^{j}\big{(}(\boldsymbol{\Sigma}^{\mathbf{x}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})^{j}\big{)}^{-1}\mathbf{G}^{\mathbf{x},j}-\mathbf{G}^{\mathbf{y},j}\right]\Delta t+O(\Delta t^{2})\in\mathbb{C}^{l\times l},$ (3.5)
$\begin{split}\mathbf{F}^{j}&:=-\mathbf{R}_{\text{{f}}}^{j}\Big{[}(\mathbf{K}^{j})^{\dagger}+\left((\mathbf{G}^{\mathbf{x},j})^{\dagger}\mathbf{K}^{j}\mathbf{R}_{\text{{f}}}^{j}(\mathbf{K}^{j})^{\dagger}-(\mathbf{R}_{\text{{f}}}^{j})^{-1}(\mathbf{H}^{j})^{\dagger}\mathbf{R}_{\text{{f}}}^{j}(\mathbf{K}^{j})^{\dagger}+(\boldsymbol{\Lambda}^{\mathbf{y},j})^{\dagger}(\mathbf{K}^{j})^{\dagger}\right)\Delta t\\ &\hskip 42.67912pt-(\boldsymbol{\Lambda}^{\mathbf{x},j})^{\dagger}\left(\big{(}(\boldsymbol{\Sigma}^{\mathbf{x}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})^{j}\big{)}^{-1}+\mathbf{K}^{j}\mathbf{R}_{\text{{f}}}^{j}(\mathbf{K}^{j})^{\dagger}\Delta t\right)\Big{]}+O(\Delta t^{2})\in\mathbb{C}^{l\times k},\end{split}$ (3.6)
where
$\begin{gathered}\mathbf{G}^{\mathbf{x},j}:=\boldsymbol{\Lambda}^{\mathbf{x},j}+(\boldsymbol{\Sigma}^{\mathbf{x}}\circ\boldsymbol{\Sigma}^{\mathbf{y}})^{j}(\mathbf{R}_{\text{{f}}}^{j})^{-1}\in\mathbb{C}^{k\times l},\quad\mathbf{G}^{\mathbf{y},j}:=\boldsymbol{\Lambda}^{\mathbf{y},j}+(\boldsymbol{\Sigma}^{\mathbf{y}}\circ\boldsymbol{\Sigma}^{\mathbf{y}})^{j}(\mathbf{R}_{\text{{f}}}^{j})^{-1}\in\mathbb{C}^{l\times l},\\ \mathbf{H}^{j}:=(\mathbf{R}_{\text{{f}}})^{-1}\left(\boldsymbol{\Lambda}^{\mathbf{y},j}\mathbf{R}_{\text{f}}^{j}+\mathbf{R}_{\text{f}}^{j}(\boldsymbol{\Lambda}^{\mathbf{y},j})^{\dagger}+(\boldsymbol{\Sigma}^{\mathbf{y}}\circ\boldsymbol{\Sigma}^{\mathbf{y}})^{j}\right)\in\mathbb{C}^{l\times l},\quad\mathbf{K}^{j}:=\big{(}(\boldsymbol{\Sigma}^{\mathbf{x}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})^{j}\big{)}^{-1}\mathbf{G}^{\mathbf{x},j}\in\mathbb{C}^{k\times l}.\end{gathered}$ (3.7)
We also note here that in the absence of noise cross-interaction, i.e., $\boldsymbol{\Sigma}^{\mathbf{x}}\circ\boldsymbol{\Sigma}^{\mathbf{y}}\equiv\mathbf{0}_{k\times l}$, (3.5) and (3.6) simplify significantly, where up to leading-order $O(\Delta t)$ their expressions reduce down to [68]:
$\mathbf{E}^{j}$ $:=\mathbf{I}_{l\times l}-\mathbf{G}^{\mathbf{y},j}\Delta t+O(\Delta t^{2}),$ (3.8)
$\mathbf{F}^{j}$ $:=\mathbf{G}^{\mathbf{y},j}\mathbf{R}_{\text{f}}^{j}(\boldsymbol{\Lambda}^{\mathbf{x},j})^{\dagger}\big{(}(\boldsymbol{\Sigma}^{\mathbf{x}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})^{j}\big{)}^{-1}\Delta t+O(\Delta t^{2}).$ (3.9)
Using (3.3)–(3.4), the following theorem outlines the procedure for obtaining an optimal online forward-in-time discrete smoother state estimate for the posterior Gaussian statistics. The proof of this result is provided in Appendix B.
Theorem 3.1 (Optimal Online Forward-In-Time Discrete Smoother).
Let $(\mathbf{x}(t),\mathbf{y}(t))$ satisfy (2.1)–(2.2) and assume the validity of the assumptions in Theorem 2.3 and all regularity conditions outlined in Appendix A. Suppose now the observational data for the observed process, $\mathbf{x}^{0},\mathbf{x}^{1},\mathbf{x}^{2},\ldots$, are given sequentially. When a new observation, denoted by $\mathbf{x}^{n}$ for $n\in\mathbb{N}$, becomes available, it is utilized to update all the existing optimal state estimates at time instants $t_{j}$ for $10\leq j\leq n-$ 1 and it thens provides a new state estimate at $j=n$. The discrete smoother posterior distribution $\mathbb{P}\big{(}\mathbf{y}^{j}\big{|}\mathbf{x}^{s},0\leq s\leq n\big{)}$ is conditionally Gaussian,
$\mathbb{P}\big{(}\mathbf{y}^{j}\big{|}\mathbf{x}^{s},0\leq s\leq n\big{)}\overset{{\,\rm d}}{\sim}\mathcal{N}_{l}(\boldsymbol{\mu}_{\text{s}}^{j,n},\mathbf{R}_{\text{s}}^{j,n}),$ (3.10)
and the conditional mean $\boldsymbol{\mu}_{\text{s}}^{j,n}$ and conditional covariance $\mathbf{R}_{\text{s}}^{j,n}$ for $n\in\mathbb{N}$ and $10\leq j\leq n-$ 1 satisfy the following recursive backward difference equations:
$\boldsymbol{\mu}_{\text{s}}^{j,n}$ $=\boldsymbol{\mu}_{\text{s}}^{j,n-1}+\mathbf{D}^{j,n-2}\left(\boldsymbol{\mu}_{\text{s}}^{n-1,n}-\boldsymbol{\mu}_{\text{f}}^{n-1}\right),$ (3.11)
$\mathbf{R}_{\text{s}}^{j,n}$ $=\mathbf{R}_{\text{s}}^{j,n-1}+\mathbf{D}^{j,n-2}\left(\mathbf{R}_{\text{s}}^{n-1,n}-\mathbf{R}_{\text{f}}^{n-1}\right)(\mathbf{D}^{j,n-1})^{\dagger},$ (3.12)
where the update matrix $\mathbf{D}^{j,n-1}$, or $\mathbf{D}^{j,n-2}$ after the trivial reindexing $n-1\leadsto n-2$ in the following (without loss of generality), is defined in a forward-in-time fashion as
[…]
where $\mathbf{E}^{j}$ is given up to leading-order in (3.5). For $n\in\mathbb{N}$ we have
$\boldsymbol{\mu}_{\text{s}}^{n-1,n}$ $=\mathbf{E}^{n-1}\boldsymbol{\mu}_{\text{f}}^{n}+\mathbf{b}^{n-1},$ (3.14)
$\mathbf{R}_{\text{s}}^{n-1,n}$ $=\mathbf{E}^{n-1}\mathbf{R}_{\text{f}}^{n}(\mathbf{E}^{n-1})^{\dagger}+\mathbf{P}^{n-1}_{n},$ (3.15)
with the $\mathbf{b}^{n-1}$ and $\mathbf{P}^{n-1}_{n}$ auxiliary residual terms being defined by
$\begin{split}\mathbf{b}^{n-1}&:=\boldsymbol{\mu}_{\text{f}}^{n-1}-\mathbf{E}^{n-1}\left((\mathbf{I}_{l\times l}+\boldsymbol{\Lambda}^{\mathbf{y},n-1}\Delta t)\boldsymbol{\mu}_{\text{{f}}}^{n-1}+\mathbf{f}^{\mathbf{y},n-1}\Delta t\right)\\ &\hskip 42.67912pt+\mathbf{F}^{n-1}\left(\mathbf{x}^{n}-\mathbf{x}^{n-1}-(\boldsymbol{\Lambda}^{\mathbf{x},n-1}\boldsymbol{\mu}_{\text{{f}}}^{n-1}+\mathbf{f}^{\mathbf{x},n-1})\Delta t\right),\end{split}$ (3.16)
$\mathbf{P}^{n-1}_{n}$ $:=\mathbf{R}_{\text{f}}^{n-1}-\mathbf{E}^{n-1}(\mathbf{I}_{l\times l}+\boldsymbol{\Lambda}^{\mathbf{y},n-1}\Delta t)\mathbf{R}_{\text{f}}^{n-1}-\mathbf{F}^{n-1}\boldsymbol{\Lambda}^{\mathbf{x},n-1}\mathbf{R}_{\text{f}}^{n-1}\Delta t,$ (3.17)
for $\mathbf{E}^{n}$ and $\mathbf{F}^{n}$ given up to leading-order in (3.5) and (3.6), respectively. For $j=n$, we have by definition that $\boldsymbol{\mu}_{\text{s}}^{n,n}=\boldsymbol{\mu}_{\text{f}}^{n}$ and $\mathbf{R}_{\text{s}}^{n,n}=\mathbf{R}_{\text{f}}^{n}$, since the smoother and filter posterior Gaussian statistics coincide at the end point per Theorem 2.3.
Based on this theorem, the algorithm associated with the online forward-in-time discrete smoother update equations (3.11)–(3.12) is outlined in Algorithm 1 in the form of pseudocode. Notice that in Algorithm 1 we use a compact expression for the update matrix $\mathbf{D}^{j,n-2}$,
$\mathbf{D}^{j,n-2}=\overset{{\curvearrowright}}{\prod^{n-2}_{i=j}}\mathbf{E}^{i}=\mathbf{E}^{j}\mathbf{E}^{j+1}\cdots\mathbf{E}^{n-2},$ (3.18)
which is equivalent to (3.13), with the details of this equivalence being shown in Appendix B.
Data: $\mathbf{x}^{0}$, $\boldsymbol{\mu}_{\text{f}}^{0}=\boldsymbol{\mu}_{\text{f}}(0)$, $\mathbf{R}_{\text{f}}^{0}=\mathbf{R}_{\text{f}}(0)$, $\Delta t$
Result: Discrete Smoother Gaussian Statistics $\left\{\boldsymbol{\mu}_{\text{s}}^{j,n}\right\}_{0\leq j\leq n}$ and $\left\{\mathbf{R}_{\text{s}}^{j,n}\right\}_{0\leq j\leq n}$ for $n\in\mathbb{N}$
for $n\in\mathbb{N}$ do
       Receive new observation $\mathbf{x}^{n}$ after time $\Delta t$;
       Compute $\boldsymbol{\mu}_{\text{s}}^{n,n}=\boldsymbol{\mu}_{\text{f}}^{n}$ through (2.3):
         $\boldsymbol{\mu}_{\text{f}}^{n}\leftarrow\boldsymbol{\mu}_{\text{f}}^{n-1}+\left(\boldsymbol{\Lambda}^{\mathbf{y},n-1}\boldsymbol{\mu}_{\text{f}}^{n-1}+\mathbf{f}^{\mathbf{y},n-1}\right)\Delta t+\left(\mathbf{R}_{\text{f}}^{n-1}(\boldsymbol{\Lambda}^{\mathbf{x},n-1})^{\dagger}+(\boldsymbol{\Sigma}^{\mathbf{y}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})^{n-1}\right)$
                  ×$\times\big{(}(\boldsymbol{\Sigma}^{\mathbf{x}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})^{n-1}\big{)}^{-1}\left(\mathbf{x}^{n}-\mathbf{x}^{n-1}-\left(\boldsymbol{\Lambda}^{\mathbf{x},n-1}\boldsymbol{\mu}_{\text{f}}^{n-1}+\mathbf{f}^{\mathbf{x},n-1}\right)\Delta t\right)$;
       Compute $\mathbf{R}_{\text{s}}^{n,n}=\mathbf{R}_{\text{f}}^{n}$ through (2.4):
         $\mathbf{R}_{\text{f}}^{n}\leftarrow\mathbf{R}_{\text{f}}^{n-1}+\Big{(}\boldsymbol{\Lambda}^{\mathbf{y},n-1}\mathbf{R}_{\text{f}}^{n-1}+\mathbf{R}_{\text{f}}^{n-1}(\boldsymbol{\Lambda}^{\mathbf{y},n-1})^{\dagger}+(\boldsymbol{\Sigma}^{\mathbf{y}}\circ\boldsymbol{\Sigma}^{\mathbf{y}})^{n-1}$
                   $-\big{(}\mathbf{R}_{\text{f}}^{n-1}(\boldsymbol{\Lambda}^{\mathbf{x},n-1})^{\dagger}+(\boldsymbol{\Sigma}^{\mathbf{y}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})^{n-1}\big{)}\big{(}(\boldsymbol{\Sigma}^{\mathbf{x}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})^{n-1}\big{)}^{-1}\big{(}\boldsymbol{\Lambda}^{\mathbf{x},n-1}\mathbf{R}_{\text{f}}^{n-1}+(\boldsymbol{\Sigma}^{\mathbf{x}}\circ\boldsymbol{\Sigma}^{\mathbf{y}})^{n-1}\big{)}\Big{)}\Delta t$;
       Compute $\mathbf{E}^{n-1}$, $\mathbf{b}^{n-1}$, and $\mathbf{P}^{n-1}_{n}$ through (3.5), (3.16), and (B), respectively:
         $\mathbf{E}^{n-1}\leftarrow\mathbf{I}_{l\times l}+\Big{(}(\boldsymbol{\Sigma}^{\mathbf{y}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})^{n-1}\big{(}(\boldsymbol{\Sigma}^{\mathbf{x}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})^{n-1}\big{)}^{-1}\mathbf{G}^{\mathbf{x},n-1}-\mathbf{G}^{\mathbf{y},n-1}\Big{)}\Delta t$;
         $\mathbf{b}^{n-1}\leftarrow\boldsymbol{\mu}_{\text{f}}^{n-1}-\mathbf{E}^{n-1}\left((\mathbf{I}_{l\times l}+\boldsymbol{\Lambda}^{\mathbf{y},n-1}\Delta t)\boldsymbol{\mu}_{\text{{f}}}^{n-1}+\mathbf{f}^{\mathbf{y},n-1}\Delta t\right)$
                     $+\mathbf{F}^{n-1}\left(\mathbf{x}^{n}-\mathbf{x}^{n-1}-(\boldsymbol{\Lambda}^{\mathbf{x},n-1}\boldsymbol{\mu}_{\text{{f}}}^{n-1}+\mathbf{f}^{\mathbf{x},n-1})\Delta t\right)$;
         $\mathbf{P}_{n}^{n-1}\leftarrow\mathbf{R}_{\text{f}}^{n-1}-\mathbf{E}^{n-1}(\mathbf{I}_{l\times l}+\boldsymbol{\Lambda}^{\mathbf{y},n-1}\Delta t)\mathbf{R}_{\text{f}}^{n-1}-\mathbf{F}^{n-1}\boldsymbol{\Lambda}^{\mathbf{x},n-1}\mathbf{R}_{\text{f}}^{n-1}\Delta t$;
      Compute $\boldsymbol{\mu}_{\text{s}}^{n-1,n}$ and $\mathbf{R}_{\text{s}}^{n-1,n}$ through (3.14) and (3.15), respectively:
         $\boldsymbol{\mu}_{\text{s}}^{n-1,n}=\mathbf{E}^{n-1}\boldsymbol{\mu}_{\text{f}}^{n}+\mathbf{b}^{n-1}$;
         $\mathbf{R}_{\text{s}}^{n-1,n}=\mathbf{E}^{n-1}\mathbf{R}_{\text{f}}^{n}(\mathbf{E}^{n-1})^{\dagger}+\mathbf{P}^{n-1}_{n}$;
      for $j=0:n-1$ do
             $\mathbf{D}^{j,n-2}\leftarrow\overset{{\curvearrowright}}{\prod^{n-2}_{i=j}}\mathbf{E}^{i}$;
             $\boldsymbol{\mu}_{\text{s}}^{j,n}\leftarrow\boldsymbol{\mu}_{\text{s}}^{j,n-1}+\mathbf{D}^{j,n-2}\big{(}\boldsymbol{\mu}_{\text{s}}^{n-1,n}-\boldsymbol{\mu}_{\text{f}}^{n-1}\big{)}$;
             $\mathbf{R}_{\text{s}}^{j,n}=\mathbf{R}_{\text{s}}^{j,n-1}+\mathbf{D}^{j,n-2}\big{(}\mathbf{R}_{\text{s}}^{n-1,n}-\mathbf{R}_{\text{f}}^{n-1}\big{)}(\mathbf{D}^{j,n-2})^{\dagger}$;
       end for
end for
Algorithm 1 Optimal Online Forward-In-Time Discrete Smoother
3.2 Understanding How the Online Forward-In-Time Discrete Smoother Works
Figure 3.1 presents a schematic diagram that intuitively explains how the online smoother operates using forward filtering and backward smoothing estimates. Although the figure focuses on the online discrete smoother mean for brevity, the same logic (as well as subsequent analysi) applies to the smoother covariance. In the diagram, the first superscript index corresponds to rows, while the second corresponds to columns. The discussion centers on the last red dashed box (last column), representing the discrete smoother estimates at the current observation (the $n$-th observation) over all time instants $j=0,\ldots,n$. Each column (fixed second index) illustrates the backward smoothing process for a given batch realization $\{\mathbf{x}^{0},\ldots,\mathbf{x}^{n}\}$, involving a forward pass to compute the filter estimate (Theorem 2.2) and a backward pass to calculate the smoother estimate (Theorem 2.3), indicated by the red arrows. Each row (fixed first index) shows how the discrete smoother estimates at a fixed time $t_{j}$ are updated as new observations arrive sequentially, using the online forward-in-time smoother algorithm. Information flows onto $\boldsymbol{\mu}_{\text{s}}^{j,n}$ from the filter estimates $\boldsymbol{\mu}_{\text{f}}^{j}$ and $\boldsymbol{\mu}_{\text{f}}^{n}$. The former, $\boldsymbol{\mu}_{\text{f}}^{j}$, influences it through forward online-smoothing in its row, transitioning towards $\boldsymbol{\mu}_{\text{s}}^{j,n-1}$ and to the next columns using the online forward-in-time smoother (3.11). Through $\boldsymbol{\mu}_{\text{s}}^{j,n-1}$, the information in $\boldsymbol{\mu}_{\text{f}}^{n-1}$ is also captured via the classical backward smoother in its column. The latter, $\boldsymbol{\mu}_{\text{f}}^{n}$, affects it via the backward smoother (3.3) in its column, implicitly through $\boldsymbol{\mu}_{\text{s}}^{n-1,n}$ in (3.11). These observations via Figure 3.1 are explicitly and mathematically stated by the fact that the amount of updated information incurred at time step $t_{j}$, due to the new observation $\mathbf{x}^{n}$, is proportional to the update at $t_{n-1}$ via the matrix $\mathbf{D}^{j,n-2}$, since
$\boldsymbol{\mu}_{\text{s}}^{j,n}-\boldsymbol{\mu}_{\text{s}}^{j,n-1}=\mathbf{D}^{j,n-2}\left(\boldsymbol{\mu}_{\text{s}}^{n-1,n}-\boldsymbol{\mu}_{\text{f}}^{n-1}\right)\quad\text{and}\quad\boldsymbol{\mu}_{\text{f}}^{n-1}=\boldsymbol{\mu}_{\text{s}}^{n-1,n-1}.$
Essentially, $\boldsymbol{\mu}_{\text{s}}^{n-1,n}-\boldsymbol{\mu}_{\text{f}}^{n-1}=\boldsymbol{\mu}_{\text{s}}^{n-1,n}-\boldsymbol{\mu}_{\text{s}}^{n-1,n-1}$ amounts to the innovation in the mean stemming from the new observational data, which is then being weighted by the optimal gain tensor $\mathbf{D}^{j,n-2}$.
Figure 3.1: Schematic diagram of how the online discrete smoother update works. Only the smoother mean $\boldsymbol{\mu}_{\text{s}}^{\cdot,\cdot}$ is depicted; the same diagram applies to the smoother covariance $\mathbf{R}_{\text{s}}^{\cdot,\cdot}$ as well, without loss of generality.
It becomes apparent from (3.11)–(3.12) that to understand how the online smoother update works, as well as to understand how information dissipates through it, we need to study the mathematical properties of the update tensor given in (3.13), or compactly in (3.18), which in turn implies the thorough study of the auxiliary operator
$\mathbf{E}^{j}=\mathbf{I}_{l\times l}-\Big{(}\mathbf{G}^{\mathbf{y},j}-(\boldsymbol{\Sigma}^{\mathbf{y}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})^{j}\big{(}(\boldsymbol{\Sigma}^{\mathbf{x}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})^{j}\big{)}^{-1}\mathbf{G}^{\mathbf{x},j}\Big{)}\Delta t+O(\Delta t^{2}),$
for $j=0,1,\ldots,n-2$. Based on the definitions of (2.6)–(2.7), it is easy to see that
$\mathbf{E}^{j}=\mathbf{I}_{l\times l}-\Big{(}\mathbf{A}^{j}+\mathbf{B}^{j}(\mathbf{R}_{\text{f}}^{j})^{-1}\Big{)}\Delta t+O(\Delta t^{2})=\mathbf{I}_{l\times l}-\Big{(}\mathbf{A}^{j}\mathbf{R}_{\text{f}}^{j}+\mathbf{B}^{j}\Big{)}(\mathbf{R}_{\text{f}}^{j})^{-1}\Delta t+O(\Delta t^{2}).$
As such, it is immediate by this form of $\mathbf{E}^{j}$, that if $\Delta t$ is sufficiently small and
$\mathbf{0}_{l\times l}\prec\mathbf{G}^{\mathbf{y},j}-(\boldsymbol{\Sigma}^{\mathbf{y}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})^{j}\big{(}(\boldsymbol{\Sigma}^{\mathbf{x}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})^{j}\big{)}^{-1}\mathbf{G}^{\mathbf{x},j}=\mathbf{A}^{j}+\mathbf{B}^{j}(\mathbf{R}_{\text{f}}^{j})^{-1},$ (3.19)
where $\prec$” is to be understood in the Loewner partial ordering sense over the convex cone of nonnegative-definite matrices, then by definition the spectral radius of $\mathbf{E}^{j}$ is less than $1$ uniformly over $j$,
$\varrho(\mathbf{E}^{j})<1,\ \forall j=0,1,\ldots,n-1.$ (3.20)
It can be shown that (3.19) is actually a necessary condition as to establish the mean-square stability of the smoother Gaussian statistics [77, 78, 79, 80, 81, 82], as can be clearly seen by the damping in the backward random differential equation that the smoother covariance satisfies in (2.9) of Theorem 2.3, with its discrete counterpart given in (3.4), and where the same damping coefficient also appears in the smoother mean’s evolution equation, since after some brief algebraic manipulations we have
$\overleftarrow{{\,\rm d}\boldsymbol{\mu}_{\text{{s}}}}=\left(-(\mathbf{A}+\mathbf{B}\mathbf{R}_{\text{{f}}}^{-1})\boldsymbol{\mu}_{\text{{s}}}+\mathbf{f}^{\mathbf{y}}-\mathbf{B}\mathbf{R}_{\text{{f}}}^{-1}\boldsymbol{\mu}_{\text{{f}}}\right){\,\rm d}t+(\boldsymbol{\Sigma}^{\mathbf{y}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})(\boldsymbol{\Sigma}^{\mathbf{x}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})^{-1}\left(\overleftarrow{{\,\rm d}\mathbf{x}}+\mathbf{f}^{\mathbf{x}}{\,\rm d}t\right),$
with its discrete counterpart given in (3.3).
Note that the filter covariance matrix $\mathbf{R}_{\text{f}}^{j}$ reflects uncertainties in both the observed and unobserved dynamics, as it is conditioned on observations. If the observed variables have much less uncertainty than the unobserved process, the eigenvalues of $(\boldsymbol{\Sigma}^{\mathbf{y}}\circ\boldsymbol{\Sigma}^{\mathbf{y}})^{j}$ will be much larger than those of $\mathbf{R}_{\text{f}}^{j}$, as suggested by discretizing the random Riccati equation (2.4), since the quadratic term is weighted by $\boldsymbol{\Gamma}:=(\boldsymbol{\Lambda}^{\mathbf{x}})^{\dagger}(\boldsymbol{\Sigma}^{\mathbf{x}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})^{-1}\boldsymbol{\Lambda}^{\mathbf{x}}$ which is being taken out from the nonnegative-definite stochastic forcing $\boldsymbol{\Sigma}^{\mathbf{y}}\circ\boldsymbol{\Sigma}^{\mathbf{y}}$. Consequently, by this reasoning and (3.5), the spectrum of $\mathbf{E}^{j}$ should concentrate near the origin and within the unit disk in the complex plane. In this scenario, observations become highly informative. They provide sufficient information to update the estimated state over a short time interval. In contrast, their influence on earlier times decays exponentially, which reflects the exponential mean-square stability of the smoother covariance equations. Numerical examples, such as the nonlinear dyad-interaction model in Section 4.1, show that even with intermittent instabilities, extreme events, and noise cross-interactions, the spectral radius of the update matrix $\mathbf{D}^{j,n-2}$ exhibits exponential decay backwards in time, with the rate increasing under the observance of extreme events in the observed time series; observing the intermittent extreme events increases the signal-to-noise ratio, thus reducing the observational uncertainty. This also leads to larger required lag values in the adaptive-lag strategy for the optimal online forward-in-time smoother discussed in Section 3.4.2, as to effectively resolve the intermittent instabilities at their onset.
To make things more concrete, we show that in the case of a simple two-dimensional linear Gaussian model, the value of $\mathbf{E}^{j}\equiv E^{j}$ (which in this case is just a scalar), is necessarily in $(-1,1)$, thus concretely establishing (3.20) and by extension, due to the one-dimensional observable and unobserved state spaces, leads to $\mathbf{D}^{j,n-2}\equiv D^{j,n-2}\in(-1,1)$ for this specific example. Consider the following two-dimensional linear Gaussian model:
${\,\rm d}x$ $=\left(a_{11}x+a_{12}y\right)+\sigma_{1}{\,\rm d}W_{1},$
${\,\rm d}y$ $=\left(a_{21}x+a_{22}y\right)+\sigma_{2}{\,\rm d}W_{2},$
where $a_{11}$, $a_{12}$, $a_{21}$, $a_{22}\in\mathbb{R}$, $\sigma_{1}>0$, and $\sigma_{2}>0$ are constants, and $x$ and $y$ are the observed and unobserved variables, respectively. Note that we assume the absence of noise cross-interaction. Here, due to the linearity of the system, $a_{22}$ needs to be negative in order to guarantee the existence of the statistical equilibrium state of the coupled system when $y$ is assumed to be the unobservable (this also becomes apparent from (3.21) later on) [75, 83, 84, 78]. Since the coefficients are constants, the equilibrium solution for the filter covariance, $R_{\text{f}}^{\text{eq}}$, can thus be solved via the following steady state Riccati equation (see (2.4)) [42, 47, 75]:
$2a_{22}R_{\text{f}}^{\text{eq}}+\sigma_{2}^{2}=\frac{(R_{\text{f}}^{\text{eq}}a_{12})^{2}}{\sigma_{1}^{2}},$ (3.21)
which after solving for $R_{\text{f}}^{\text{eq}}$ gives,
$R_{\text{f}}^{\text{eq}}=\frac{a_{22}\sigma_{1}^{2}+\sigma_{1}\sqrt{\Psi}}{a_{12}^{2}},$ (3.22)
where $\Psi=a_{22}^{2}\sigma_{1}^{2}+a_{12}^{2}\sigma_{2}^{2}$. Observe now that by plugging-in (3.22) into the expression found in (3.19),
$\mathbf{G}^{\mathbf{y}}-(\boldsymbol{\Sigma}^{\mathbf{y}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})(\boldsymbol{\Sigma}^{\mathbf{x}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})^{-1}\mathbf{G}^{\mathbf{x}}\equiv G^{y},$
which is stationary in time and simplifies down to $G^{y}$ because there is no noise cross-interaction, then we retrieve
$G^{y}$ $=a_{22}+\frac{\sigma_{2}^{2}}{R_{\text{f}}^{\text{eq}}}=\frac{1}{R_{\text{f}}^{\text{eq}}}\left(a_{22}\frac{a_{22}\sigma_{1}^{2}+\sigma_{1}\sqrt{\Psi}}{a_{12}^{2}}+\sigma_{2}^{2}\right)$
$=\frac{1}{R_{\text{f}}^{\text{eq}}}\left(\frac{a_{22}^{2}\sigma_{1}^{2}+a_{22}\sigma_{1}\sqrt{\Psi}+a_{12}^{2}\sigma_{2}^{2}}{a_{12}^{2}}\right)=\frac{\sqrt{\Psi}}{a_{12}^{2}R_{\text{f}}^{\text{eq}}}\left(\sqrt{\Psi}+a_{22}\sigma_{1}\right)>0,$
where we have used the definition of $\Psi$, the fact that $\sqrt{\Psi}+a_{22}\sigma_{1}=\sqrt{a_{22}^{2}\sigma_{1}^{2}+a_{12}^{2}\sigma_{2}^{2}}+a_{22}\sigma_{1}>0$, and that $R_{\text{f}}^{\text{eq}}>0$. This validates condition (3.19), and so the auxiliary constant $E^{j}$ is always within the interval $(-1,1)$, meaning (3.20) is satisfied. By extension, since everything is just a scalar, this implies $\mathbf{D}^{j,n-2}\equiv D^{j,n-2}\in(-1,1)$ because of (3.18) (see Appendix B).
In general, except in the trivial cases where the unobserved state space is one-dimensional, the condition (3.20) on the constituents of the product defining the update matrix $\mathbf{D}^{j,n-2}$ in (3.18) do not guarantee that the update tensor $\mathbf{D}^{j,n-2}$ itself will also have a spectral radius less than $1$. This is because, generally, the spectral radius of a product is not dominated by the product of the associated spectral radii unless rather restrictive conditions are applied, with sufficient conditions including simultaneous triangularization of the factors over $\mathbb{C}$ [85] (e.g., the factors commute, as a consequence of Gelfand’s theorem) or the factors being radial matrices [86]. Consequently, predicting how each observation will affect state estimation through online smoother updates can be complex and highly dependent on the specific structure of the CGNS. This variability underscores the need for a simple numerical procedure to assess the impact region of each new observation, as described by the recursive backward difference equations for the discrete smoother posterior Gaussian statistics in (3.11)–(3.12), which will in turn aid in constructing an effective adaptive lag strategy for the online smoother (see Section 3.4). Nevertheless, even in the ideal setting where control on the spectrum of $\mathbf{D}^{j,n-2}$ can be established, an online smoother which defines its adaptive lag based on the eigenvalues of the update matrix will incur computational costs in the order of $O(l^{3})$, which for high-dimensional systems is inexcusable.
As the system dimension and observational period length increase, the computational overhead for updating the online smoothing procedure becomes significant, particularly due to the storage of auxiliary matrices like $\mathbf{E}^{j}$, which are essential for calculating the update matrix $\mathbf{D}^{j,n-2}$, as well as the filter and smoother covariance matrices. Additionally, constant recalculation of the covariance tensor can lead to $O(l^{3})$ computations per update, especially when $k\approx l$. Given that the impact of new observations on past states decays exponentially (linked to the aforementioned spectral properties of $\mathbf{D}^{j,n-2}$ and $\mathbf{E}^{j}$), it is advantageous to develop a fixed- or adaptive-lag online smoother [17, 27, 43, 44]. In such a way, the impact region of each observation is either finite and predetermined or computable based on the dynamical and statistical properties of the underlying system.
3.3 Fixed-Lag Strategy for the Online Smoother
The fixed-lag smoother is the simplest method for conserving computational and storage resources, as it assumes a uniformly predetermined impact region for each new observation. Setting an a priori impact region incurs no additional computational cost. However, this approach has significant drawbacks for most turbulent and complex dynamical systems: it may assign an unnecessarily long fixed lag, wasting storage with diminishing returns in state estimation, or a short fixed lag that introduces substantial biases in the estimated state. Nevertheless, it serves as a building block of the adaptive lag smoother.
The fixed-lag methodology is defined as follows: we assume a predetermined impact region size for each new observation, denoted as lag $L\in\{0,\ldots,n\}$ in time steps (edge cases $L=0$ and $L=n$ are discussed later). All subsequent measurements can be expressed in simulation time units by multiplying by $\Delta t$. We also assume that $L$ is uniform across $n$, the number of observations. The fixed-lag online smoother Gaussian statistics are then given by the following formulas:
$\boldsymbol{\mu}_{\text{s}}^{j,n}$ :=$:=\begin{cases}\boldsymbol{\mu}_{\text{s}}^{j,n-1},&\text{ for }0\leq j\leq n-1-L\\ \boldsymbol{\mu}_{\text{s}}^{j,n-1}+\mathbf{D}^{j,n-2}\big{(}\boldsymbol{\mu}_{\text{s}}^{n-1,n}-\boldsymbol{\mu}_{\text{f}}^{n-1}\big{)},&\text{ for }n-L\leq j\leq n-1,\end{cases}$ (3.23)
$\mathbf{R}_{\text{s}}^{j,n}$ :=$:=\begin{cases}\mathbf{R}_{\text{s}}^{j,n-1},&\text{ for }0\leq j\leq n-1-L\\ \mathbf{R}_{\text{s}}^{j,n-1}+\mathbf{D}^{j,n-2}\big{(}\mathbf{R}_{\text{s}}^{n-1,n}-\mathbf{R}_{\text{f}}^{n-1}\big{)}(\mathbf{D}^{j,n-2})^{\dagger},&\text{ for }n-L\leq j\leq n-1,\end{cases}$ (3.24)
with the details for the online discrete smoother already outlined in Algorithm 1 of Section 3.1. For $j=n$, as already discussed, we always have $\boldsymbol{\mu}_{\text{s}}^{n,n}=\boldsymbol{\mu}_{\text{f}}^{n}$ and $\mathbf{R}_{\text{s}}^{n,n}=\mathbf{R}_{\text{f}}^{n}$. As such, we have for $\forall n\in\mathbb{N}$ and $j=0,\ldots,n-1$, that $\boldsymbol{\mu}_{\text{s}}^{j,n}=\boldsymbol{\mu}_{\text{s}}^{j,n}(L)$ and $\mathbf{R}_{\text{s}}^{j,n}=\mathbf{R}_{\text{s}}^{j,n}(L)$, where this explicitly denotes the dependence on the predetermined fixed lag value of $L$.
We mention here that the edge cases $L=0$ and $L=n$ actually recover the optimal filter and (offline) smoother state estimations, respectively, under the consideration that we consider the $n$ observations to define a complete time series for the observable variables. It is not hard to see why this is the case. Observe that when $L=0$, then we only work with the first branch of the definitions for the fixed-lag smoother Gaussian statistics in (3.23) and (3.24), and as such with each new observation we just carry over the estimated state from the corresponding time instant $t_{j}$ we did at the previous iteration, i.e., the one from the last observation. As such, since $\boldsymbol{\mu}_{\text{s}}^{n,n}=\boldsymbol{\mu}_{\text{f}}^{n}$ and $\mathbf{R}_{\text{s}}^{n,n}=\mathbf{R}_{\text{f}}^{n}$, for every $n\in\mathbb{N}_{0}$, then it is immediate that when $L=0$, then with each new observation we just carry-on with the forward-pass and calculate the optimal filter Gaussian statistics at the new observation and then carry over all the previous state estimates. But since at each point we are calculating the filter state, then we are essentially just doing a forward-pass as the observations come in, and as such we only recover the filter posterior Gaussian statistics. On the other hand, when $L=n$, then we only work with the second branch of the definitions for the fixed-lag smoother Gaussian statistics in (3.23) and (3.24), which means that with each new observation we do the forward-pass and obtain the new filter estimate at the end-point ($j=n$), but using this new state estimation we do a full backward-pass update and obtain the smoother estimates at each and every time instant, when now considering this new observation. As such, we are essentially doing a full forward- and backward-pass with the arrival of each observation, which means we are recovering with each new observational data the offline smoother posterior Gaussian statistics, fully, for this complete observational period.
3.4 Online Smoother with an Adaptive Lag Determined Using Information Theory
The adaptive-lag method for defining optimal online smoother state estimates seeks to dynamically determine the impact region of each new observation while minimizing storage costs. To achieve a consistent strategy for calculating the required lag for each observation, we leverage fundamental tools from information theory and the conditional Gaussian structure of the CGNS framework. This approach enables the optimally estimated state from the online forward-in-time discrete smoother to more effectively capture intermittency, extreme events, and nonlinear dynamics that influence the transient behavior of complex and turbulent systems. In the following subsections, we first introduce key concepts from information theory, which contribute to the development of the adaptive-lag strategy for the online smoother.
3.4.1 Information Theory Fundamentals: Relative Entropy and the Signal-Dispersion Decomposition
Due to the turbulent nature, new observational data influence the estimated states through the online smoother update only within a finite time interval, with this impact exhibiting exponential decay over time. Therefore, it is natural to quantify the information gain from applying the optimal online forward-in-time discrete smoother (posterior) with different length of lags compared to simply carrying forward the previous state estimate (prior). By doing so, the optimal lag can be discovered. We utilize empirical information theory to develop a measure that quantifies this information gain, reflecting the additional information in the posterior distribution beyond the prior [87, 88, 89, 90, 91].
In the information-theoretic framework, a natural way to assess the information gain in the posterior when compared to the prior is through their respective PDFs by using a certain type of distance function that measures the difference between them. A general class of such functions is known as $f$-divergences [92, 93, 94].
While the general framework of $f$-divergences is rather versatile, there is a certain choice of $f$ which takes great advantage of the form of the CGNS framework, and especially of the conditional Gaussianity of the posterior distributions. Specifically, by choosing $f(z)=z\log z$, then we recover the so called Kullback-Leibler divergence or relative entropy [95, 96, 97]. In this case, we denote the divergence with $\mathcal{P}(\boldsymbol{\cdot},\boldsymbol{\cdot})$, and is simply given as [98, 99, 100]
$\mathcal{P}(p,q)=\int_{\mathbb{C}^{n}}p(\mathbf{u})\log\left(\frac{p(\mathbf{u})}{q(\mathbf{u})}\right){\,\rm d}\mathbf{u},$ (3.25)
where the integration with respect to the state vector is to be understood in the Lebesgue sense and is essentially taken over the common support of the two densities. For our framework, we read $\mathcal{P}(p,q)$ as the information gain of $p$ from $q$, as in the information that has been gained by using the true density $p$ in lieu of the approximation $q$. The relative entropy enjoys the lucrative properties of nonnegativity and of invariancy under general nonlinear changes of state variables [101]. These properties, among others, allow the relative entropy to be widely utilized in assessing the information gain and quantifying model error, model sensitivity, predictability, and statistical response for stochastic CTNDSs [102, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112].
The relative entropy ensures that the extreme events related to tail probabilities are not underestimated. The ratio of the two PDFs within the logarithm quantifies the gap in the tail probability, allowing for an unbiased characterization of statistical differences. Many CTNDSs modeled through the CGNS framework generate intermittent, rare, and extreme events. Thus, using a logarithmic generator function effectively resolves tail probability events, aiding in quantifying information gain for the adaptive lag criterion, uncertainty reduction of posterior solutions, and causal inference. In addition, relative entropy benefits from this logarithmic structure, providing a simple, closed-form formula when both distributions are Gaussian. This is ideal for the online smoother update, where both the prior and posterior distributions are Gaussian, and differ only in their statistics. Consequently, this facilitates the efficient calculation of adaptive lag, reducing computational costs and storage requirements. Specifically, when both $\mathbb{P}\overset{{\,\rm d}}{\sim}\mathcal{N}_{N}(\boldsymbol{\mu}_{\text{p}},\mathbf{R}_{\text{p}})$ and $\mathbb{Q}\overset{{\,\rm d}}{\sim}\mathcal{N}_{N}(\boldsymbol{\mu}_{\text{q}},\mathbf{R}_{\text{q}})$, where $N=\mathrm{dim}(\mathbf{u})$ is the dimension of the phase space, then the relative entropy adopts an explicit and simple formula [2, 105],
$\mathcal{P}(p,q)=\frac{1}{2}(\boldsymbol{\mu}_{\text{p}}-\boldsymbol{\mu}_{\text{q}})^{\dagger}\mathbf{R}_{\text{q}}^{-1}(\boldsymbol{\mu}_{\text{p}}-\boldsymbol{\mu}_{\text{q}})+\frac{1}{2}\left(\mathrm{tr}(\mathbf{R}_{\text{p}}\mathbf{R}_{\text{q}}^{-1})-N-\log(\mathrm{det}(\mathbf{R}_{\text{p}}\mathbf{R}_{\text{q}}^{-1}))\right),$ (3.26)
where $\mathrm{tr}$ and $\mathrm{det}$ are the trace and determinant of a matrix, respectively. The first quadratic form term in (3.26) is called the signal and measures the information gain in the mean weighted by the model or approximation covariance, whereas the second term is called the dispersion and involves only the covariance ratio $\mathbf{R}_{\text{p}}\mathbf{R}_{\text{q}}^{-1}$. This is why the expression in (3.26) is known as the signal-dispersion decomposition of the relative entropy for Gaussian variables.
3.4.2 Development of the Adaptive-Lag Strategy for the Online Smoother
Here we outline the information-theoretic approach for defining the adaptive lag at the acquirement of a new observation. Let us assume that we have currently observed the $n$-th measurement, $\mathbf{x}^{n}$, which is used to update all the existing online smoother posterior Gaussian statistics at time instants $10\leq j\leq n-$ 1, and then provide a new state estimate at $j=n$. As shown in Theorem 3.1, that update happens via (3.11)–(3.12). We would like to define an adaptive lag at the $n$-observation, $L_{n}\in\{1,\ldots,n-1\}$, such that only “the most informative or impactful” of these updates are actually carried out, in other words,
$\boldsymbol{\mu}_{\text{s}}^{j,n}\equiv\boldsymbol{\mu}_{\text{s}}^{j,n}(L_{n})$ :=$:=\begin{cases}\boldsymbol{\mu}_{\text{s}}^{j,n-1},&\text{ for }0\leq j\leq n-1-L_{n}\\ \boldsymbol{\mu}_{\text{s}}^{j,n-1}+\mathbf{D}^{j,n-2}\big{(}\boldsymbol{\mu}_{\text{s}}^{n-1,n}-\boldsymbol{\mu}_{\text{f}}^{n-1}\big{)},&\text{ for }n-L_{n}\leq j\leq n-1,\end{cases}$ (3.27)
$\mathbf{R}_{\text{s}}^{j,n}\equiv\mathbf{R}_{\text{s}}^{j,n}(L_{n})$ :=$:=\begin{cases}\mathbf{R}_{\text{s}}^{j,n-1},&\text{ for }0\leq j\leq n-1-L_{n}\\ \mathbf{R}_{\text{s}}^{j,n-1}+\mathbf{D}^{j,n-2}\big{(}\mathbf{R}_{\text{s}}^{n-1,n}-\mathbf{R}_{\text{f}}^{n-1}\big{)}(\mathbf{D}^{j,n-2})^{\dagger},&\text{ for }n-L_{n}\leq j\leq n-1,\end{cases}$ (3.28)
with the details for the online discrete smoother already outlined in Algorithm 1 of Section 3.1. For $j=n$, as already discussed, we always have $\boldsymbol{\mu}_{\text{s}}^{n,n}=\boldsymbol{\mu}_{\text{f}}^{n}$ and $\mathbf{R}_{\text{s}}^{n,n}=\mathbf{R}_{\text{f}}^{n}$.
To define the adaptive lag $L_{n}$ using the information theory principles from Section 3.4.1, we quantify the information gain from updating the state estimate at the $j$-th time step with the $n$-th observation, compared to using the estimate from the previous $(n-1)$-st observation. In the CGNS framework, two possible Gaussian distributions arise for the online smoother state estimation. The first, $p^{j,n}_{\text{updated}}$, is the optimal posterior distribution obtained through the online forward-in-time smoother, discussed in Section 3.1. The second, $p^{j,n}_{\text{lagged}}$, is the prior distribution, where no update is made and the previous state estimate at time $t_{j}$ is carried over instead. The posterior or the “updated” distribution has Gaussian statistics given by
$\boldsymbol{\mu}_{\text{s}}^{j,n-1}+\mathbf{D}^{j,n-2}\left(\boldsymbol{\mu}_{\text{s}}^{n-1,n}-\boldsymbol{\mu}_{\text{f}}^{n-1}\right)\quad\text{and}\quad\mathbf{R}_{\text{s}}^{j,n-1}+\mathbf{D}^{j,n-2}\left(\mathbf{R}_{\text{s}}^{n-1,n}-\mathbf{R}_{\text{f}}^{n-1}\right)(\mathbf{D}^{j,n-2})^{\dagger},$
for its posterior mean and covariance tensor, respectively, as indicated in (3.11)–(3.12). On the other hand, the prior or “lagged” distribution has the corresponding Gaussian statistics of
$\boldsymbol{\mu}_{\text{s}}^{j,n-1}\quad\text{and}\quad\mathbf{R}_{\text{s}}^{j,n-1},$
i.e., no update (or innovation, under the Kalman gain terminology) is being made or procured due to the newly acquired observation. This is precisely described by the first and second branches of (3.27)–(3.28), respectively.
With the posterior or “updated” conditional Gaussian distribution denoting the true distribution, while the prior or “lagged” conditional Gaussian distribution denotes the approximation due to the adaptive lag (by not carrying out the online smoother update), we can then naturally take advantage of the signal-dispersion decomposition formula for the relative entropy given in (3.26). As such, we have that the information gain or uncertainty reduction incurred by carrying out the update through the online smoother at time $t_{j}$ by using the newly obtained $n$-th observation, is equal to
$\begin{split}\mathcal{P}\left(p^{j,n}_{\text{updated}},p^{j,n}_{\text{lagged}}\right)=&\ \frac{1}{2}\left(\boldsymbol{\mu}_{\text{s}}^{n-1,n}-\boldsymbol{\mu}_{\text{f}}^{n-1}\right)^{\dagger}(\mathbf{D}^{j,n-2})^{\dagger}(\mathbf{R}_{\text{s}}^{j,n-1})^{-1}\mathbf{D}^{j,n-2}\left(\boldsymbol{\mu}_{\text{s}}^{n-1,n}-\boldsymbol{\mu}_{\text{f}}^{n-1}\right)\\ &+\frac{1}{2}\left(\mathrm{tr}(\mathbf{Q}^{j,n})-l-\mathrm{ln}(\mathrm{det}(\mathbf{Q}^{j,n}))\right),\end{split}$ (3.29)
where $\mathbf{Q}^{j,n}$ is defined as a covariance ratio matrix,
$\begin{split}\mathbf{Q}^{j,n}&:=\left(\mathbf{R}_{\text{s}}^{j,n-1}+\mathbf{D}^{j,n-2}\left(\mathbf{R}_{\text{s}}^{n-1,n}-\mathbf{R}_{\text{f}}^{n-1}\right)(\mathbf{D}^{j,n-2})^{\dagger}\right)(\mathbf{R}_{\text{s}}^{j,n-1})^{-1}\\ &=\mathbf{I}_{l\times l}+\left(\mathbf{D}^{j,n-2}\left(\mathbf{R}_{\text{s}}^{n-1,n}-\mathbf{R}_{\text{f}}^{n-1}\right)(\mathbf{D}^{j,n-2})^{\dagger}\right)(\mathbf{R}_{\text{s}}^{j,n-1})^{-1}.\end{split}$ (3.30)
The first term in (3.29) corresponds to the signal, while the latter to the dispersion, terms already introduced in Section 3.4.1. With (3.29)–(3.30) in-hand, we can now outline the procedure for determining $L_{n}$. Having a predefined upper lag bound, hereby denoted by $b\in\mathbb{N}$, we calculate
$\mathcal{P}\left(p^{j,n}_{\text{updated}},p^{j,n}_{\text{lagged}}\right)\text{for }R_{n}\leq j\leq n-1,$
where $R_{n}:=\max\{n-1-b,1\}$, thus totaling
$n-R_{n}=\begin{cases}b+1,&\text{ if }n-1-b\geq 1\\ n-1,&\text{ if }n-1-b<1,\end{cases}$
[…]
$L_{n}\equiv L_{n}(b,w,\delta):=n-1-\mathrm{max}\Big{\{}R_{n}\leq j\leq n-1:\sigma^{j,n}(w)<\delta\Big{\}}\Big{(}\in\{0,1,\ldots,b\}\Big{)},$ (3.31)
where if no such maximizer exists, then we simply set
$L_{n}:=\min\{n-1,b\}.$ (3.32)
Notice that for larger values of $\delta$ we allow for more relaxed or smaller lags, while for smaller tolerance values we push the adaptive lag values towards their predefined upper bound, $b$. Per our methodology, and as noted in (3.31), we have that $\forall n\in\mathbb{N}$, $L_{n}\in\{0,1,\ldots,b\}$.
As aforementioned, it is possible to instead use the sequence of information gains directly to define the adaptive lag,
$L_{n}\equiv L_{n}(b,\delta):=n-1-\mathrm{max}\Big{\{}R_{n}\leq j\leq n-1:\mathcal{P}\left(p^{j,n}_{\text{updated}},p^{j,n}_{\text{lagged}}\right)<\delta\Big{\}},$ (3.33)
where again if no such maximiser exists we simply use (3.32). For most cases (cf. the numerical experiments in Sections 4.1 and 4.2 where $\sigma^{j,n}$ and $\mathcal{P}\left(p^{j,n}_{\text{updated}},p^{j,n}_{\text{lagged}}\right)$ behave as $C^{n}_{1}e^{C^{n}_{2}j}$ for $R_{n}\leq j\leq n-1$ and $C^{n}_{1},C^{n}_{2}>0$) leads to the same results after a slight modification to the tolerance parameter being used (and for appropriate values of $w$ in (3.31)). In such settings, if we use $\delta=O(10^{-d})$ with the sequence of local standard deviations, for $d\in\mathbb{N}$, then we can get similar results when using the original sequence of relative entropies by using $\delta=O(10^{-d+s/2})$ where $s\in\mathbb{N}$ is defined through the time step as $\Delta t=O(10^{-s})$, since the former is approximately the first derivative of the latter with respect to the time step, as already mentioned before. The simplicity and flexibility of the adaptive lag framework in (3.31) or (3.33) further allows for the inclusion of a penalty for large lag values. This can be achieved by modifying the criterion (e.g., by using a monotonic polynomial of time as the penalization) and then choosing the largest lag for which this penalized information gain (or its penalized local standard deviation) remains small under the tolerance $\delta$, while prioritizing smaller lags.
4 Applications of the Adaptive-Lag Online Smoother
This section demonstrates the application of the online smoother in addressing three key scientific challenges: (a) state estimation and causality detection, especially in the presence of intermittency and extreme events, (b) evaluating computational and storage efficiency in a high-dimensional LDA application, and (c) developing an online parameter estimation algorithm with partial observations.
The simulations for all the numerical experiments in this work were carried out on an AMD RyzenTM 7 5800H mobile CPU of 8 cores, 16 threads, with an average clock speed of 3.2 GHz and turbo boost up to 4.4 GHz, with a TDP of 45W. In terms of memory, a two-memory-module configuration was used, with 2x16GB DDR4 RAM sticks at 3200MT/s.
4.1 Detecting Causal Evidence Using the Adaptive-Lag Online Smoother
Causal inference helps identify the mechanisms that drive the dynamics of a system. It advances more accurate predictions, better control, and improved decision-making under uncertainty [113, 114]. It also reveals the role of extreme events and intermittency in modulating the dynamical evolution of the underlying system. By clarifying cause-effect relationships, causal inference enhances our ability to model, simulate, and intervene in complex systems. This subsection demonstrates how the adaptive online smoother and the associated information gain, using a simple nonlinear dyad model, can reveal causal links by showing how time-delayed information in one process improves state estimation in another. The model generates intermittency and extreme events, and the smoother identifies how long temporal information is needed for causal detection during different phases.
The dyad model considered here is the following stochastic model with quadratic nonlinearities, multiplicative and cross-interacting noise, and physics-constraints (energy conservation in the quadratic nonlinear terms):
${\,\rm d}u(t)$ $=(-d_{u}u(t)+\gamma u(t)v(t)+F_{u}){\,\rm d}t+\sigma_{u}{\,\rm d}W_{u}(t),$ (4.1)
${\,\rm d}v(t)$ $=(-d_{v}v(t)-\gamma u(t)^{2}+F_{v}){\,\rm d}t+\sigma_{vu}u(t){\,\rm d}W_{u}(t)+\sigma_{v}{\,\rm d}W_{v}(t),$ (4.2)
where all parameters are constants, with $d_{u},d_{v}\in\mathbb{R}$ and $\gamma,\sigma_{u},\sigma_{vu},\sigma_{v}>0$, while $W_{u}$ and $W_{v}$ are two mutually independent Wiener processes. In (4.1)–(4.2), The variable $u$ represents an observed or resolved mode in the turbulent signal, interacting with the unresolved mode $v$ through quadratic nonlinearities. As $v$ influences the damping of $u$, variations in $v$ cause the time series of $u$ to display intermittency and extreme events. The system (4.1)–(4.2) fits exactly the CGNS framework in (2.1)–(2.2), with $\mathbf{x}=x:=u$ and $\mathbf{y}=y:=v$. The following parameter values are used for this numerical case study:
$d_{u}=0.5,\quad\gamma=3,\quad F_{u}=1,\quad\sigma_{u}=0.6,$
$d_{v}=0.5,\quad F_{v}=0.3,\quad\sigma_{vu}=0.8,\quad\sigma_{v}=1.$
The parameters are assigned to induce intermittent and extreme events in the observable variable $u$ through the stochastic damping generated by the unobserved variable $v$, with the goal of producing highly non-Gaussian PDFs for both observable and unobserved dynamics. These values are also chosen to address potential issues of observability [13, 115, 116, 117, 118], which is linked to causal detection using the adaptive online smoother. The coupled system in (4.1)–(4.2) loses practical observability when the observed process $u$ provides no information about the unobserved variable $v$. Intuitively, this occurs during phases where $u\approx 0$, rendering $v$ ineffective in contributing to the dynamics of $u$. A numerical integration time step of $\Delta t=0.005$ is used, which also defines the uniform rate of obtaining the observations for the online smoother, with a total simulation time of $T=60$ units.
Figure 4.1 presents various plots comparing the filter and smoother posterior estimated states at different levels. Panel (a) displays the trajectory of the observed variable $u$, while panel (b) displays that for the unobserved variable $v$. For the latter, we also include the posterior means calculated using Theorems 2.2 and 2.3, along with the first two standard deviations away from the mean, represented by accordingly colored shaded regions. The posterior estimated states overall follow the true values with relatively low uncertainty. As expected, the smoother provides a more accurate estimation during the uncertain quiescent periods of the unobserved variable. During the extreme events which are characterized by a high signal-to-noise ratio, both methods lead to smaller uncertainty. Panel (c) plots the information gain of the filter and smoother posterior distributions relative to the prior equilibrium distribution of $v$ on a logarithmically scaled y-axis. Since both the equilibrium (computed numerically using many long simulated trajectories; not shown here) and posterior distributions are Gaussian, the information gain in panel (c) is calculated using the signal-dispersion decomposition formula of relative entropy in (3.26). The temporal evolution of relative entropy reveals that the information gain from the smoother posterior Gaussian distribution is almost everywhere larger than that from the filter solution, as is expected. Notably, the information gain corresponding to the onset of the intermittent phases of $u$ shoots up due to the strong signal-to-noise ratios. Note that the information gain using the smoother is more significant than the filter at these phases and the more uncertain quiescent phases due to additional observational information. Panel (d) shows the PDF of $u$ generated by its complete signal over $t\in[0,60]$, alongside the corresponding Gaussian fit, i.e., the normal distribution with the same mean and standard deviation as the truth over the associated period. Due to the intermittent extreme events, we can discern that the induced density is highly non-Gaussian, with an apparent positive skewness and a heavy tail. Panel (e) is the same as (d) for $v$, which also includes the densities corresponding to the filter and smoother solutions. We can again see that the density corresponding to $v$ displays some non-Gaussian features. Furthermore, as is expected, the PDF of the smoother solution is better equipped to approximate that of the truth compared to the filter one by better resolving the tail behavior and first few ordered moments.
Figure 4.1: Panel (a): Trajectory of the observable variable $u$. Panel (b): True trajectory of the unobserved variable $v$ (in blue), alongside the filter (in green) and smoother (in red) posterior mean time series. Their respective first two standard deviations away from the mean state are also plotted through correspondingly colored shaded regions, with the standard deviation in this case being exactly equal to the square root of the respective posterior covariance. The threshold above which $v$ acts as anti-damping to $u$, i.e., the line $y=d_{u}/\gamma=1/6$, is also plotted in a dashed black line. Panel (c): Temporal evolution of the information gain of the filter and smoother posterior Gaussian statistics beyond the statistical attractor of the unobserved variable. This is calculated using the signal-dispersion decomposition of the relative entropy in (3.26). A logarithmic scale is used for the y-axis. Panel (d): PDF of $u$ calculated using the observations over $t\in[0,60]$ (in blue), alongside its Gaussian fit density defined by the mean and standard deviation of the signal in the same time period (in dashed black). Panel (e): Same as (d) but for the unobserved variable $v$, together with the PDFs corresponding to the filter (in green) and smoother (in red) posterior mean time series.
Figure 4.2 focuses on a shorter intermittent period ($t\in[32.5,40]$) to assess the performance of the adaptive-lag online smoother, using the LSDF sequence, $\{\sigma^{j,n}\}_{R_{n}\leq j\leq n-1}$, with the hyperparameters set to $b=400$ (which corresponds to $b\Delta t=2$ simulation time units), $\delta=10^{-6}$, and using MATLAB’s default value for $w$ in stdfilt which is $w=3$. For this case study, equivalent results can be obtained when instead the entropy sequence itself, $\left\{\mathcal{P}\left(p^{j,n}_{\text{updated}},p^{j,n}_{\text{lagged}}\right)\right\}_{R_{n}\leq j\leq n-1}$, is being used, as already discussed in Section 3.4.2, by appropriately adjusting $\delta$. Panel (a) shows the signals of $u$ and $v$ over this period, while panels (b)–(f) illustrate how the online smoother adjusts its state estimation as new observations arrive at five different time instants during this period. The online smoother is compared against the filter solution, which can also produce real-time estimated states with the sequential arrival of observations. We focus on panels (c) and (d). When using the filter solution, the estimation of $v$ before $u$ reaches the peak of the extreme event contains significant errors. This is because the growth of $v$, once it exceeds the threshold of $d_{u}/\gamma=1/6$ to become anti-damping, triggers the extreme event in $u$ (the cause). However, the extreme event (the effect) does not appear instantly, thus introducing a delay in the onset of these instabilities. This is why at $t=36.50$ (panel (c)), without seeing the future information from the development of the extreme event, the causal variable $v$ cannot be accurately recovered by the filter. By extension, due to the data from the extreme event in the observed signal simply not being available yet at the current time instant, this, in a sense, contaminates the online smoother state, which faithfully follows the filter one regardless of how long the adaptive lag is. Nevertheless, at $t=37.00$ (panel (d)), with the observed trajectory currently undergoing the effects of the instability caused by $v$, this allows the adaptive online smoother, which utilizes future information for state estimation unlike the filter one, to correct its previous estimated states based on the value of the calculated adaptive lag, at most for $b\Delta t=2$ time units into the past. This is explicitly seen by the online smoother posterior mean, which now better approximates the true signal compared to its state half a time unit ago, thus demonstrating its property to systematically correct biases when future data becomes available. This attribute is crucial to help identify the current state, therefore, outperforming the filter solution. This crucial discrepancy between the filter and online smoother estimates reveals that information in the dynamics flows from the current state of the unobserved variable (cause) to the future state of the observed one (effect), highlighting a causal relationship between these states. Panel (g) presents the adaptive lag values generated during this period by using the LSDFs of the information gain sequence outlined in (3.31). The lag in the online smoother likewise reflects the aforementioned time delay in the causal relationship, with a peak in adaptive lag values at the onset of the extreme event, i.e., the causal period. These are essential to correct past errors and effectively recover the state of the cause $v$, which triggered the extreme event. In contrast, smaller lag values are needed during the effect period, specifically during the demise of this intermittent event, which is characterized by a high signal-to-noise ratio where both the filter and smoother estimates can effectively capture the relative deterministic behavior. As such, lower adaptive lag values are required since the “lagged” distribution in (3.29) can yield a quickly decaying information gain function as $j$ decreases from $n-1$ to $R_{n}$ (or to an already small sequence of information gains for these values of $n$, i.e., below the threshold parameter $\delta$ uniformly in $j$), thus leading to small values for $L_{n}$ in (3.31). After all, the lag essentially functions as an interpolation between the past state estimates and a “full” (since $b$ limits how much into the past we can correct) backward-run smoother solution of the online smoother. As such, a close-to-zero lag value signifies that more trust is put in the statistics from the previously calculated online posterior distributions since the observations obtained during this period have tiny impact regions.
In panel (h) of Figure 4.2, the LSDF criterion sequences used to define the adaptive lag in panel (g) (see (3.31)) at the time instants showed in panels (b) through (f) are plotted, on a logarithmically scaled y-axis. Although each one is defined for $R_{n}\leq j\leq n-1$, here these are plotted on a common x-axis for ease of comparison, with the lag on the x-axis also being measured in simulation time units for clarity, i.e., the x-axis is over $[0,b\Delta t]$. The tolerance value of $\delta=10^{-6}$ is also plotted in a black dashed line. Based on (3.31), $b\Delta t=2$ minus the first time on which each curve goes below this threshold line, defines the respective adaptive lag value (i.e., for the corresponding $n$) in panel (g). We note that the minute oscillations observed are by virtue of the LSDF sequence essentially functioning as the numerical first-order derivative of the information gains. Increasing the window size $w$ or using the sequence of relative entropies instead yields equivalent curves but without these numerical artifacts. Of significance are the curves that cover the cause and effect of the extreme event around $[36.5,37.5]$, specifically the curves for $t=37.00$ (in green) and $t=38.50$ (in magenta). The former only captures its cause, so the damping in the information gain criterion is only barely observed for $j$ close to $n-1$ (i.e., close to $b\Delta t=2$ in this common and absolute x-axis). On the other hand, for the curve corresponding to $t=38.50$, with $j$ decreasing from $n-1$ down to $R_{n}$ (i.e., from $b\Delta t=2$ down to 00), we can see that there is a significant change in the exponential behavior of the information gain criterion at around $t=37.50$ which is the demise point of the extreme event in $u$, with the high signal-to-noise ratio extreme event between $[36.5,37.5]$ inducing a regime switch in the information gain criterion to a smaller exponential function with roughly the same rate, through sub-unit damping, which in turn yields the significant drop in the gained information as a function of $j$. This has the ulterior outcome of smaller adaptive lag values during this period, as shown in panel (g), compared to the other periods. For the sake of completion, observe how this behavior also starts to form for the green curve corresponding to $t=37.00$ for $j$ close to $R_{n}$ (i.e., close to 00) since it captures the demise of the effect in the previous (and milder) extreme event observed at around $t=35$. In contrast, the blue curve for $t=40.00$ is almost a straight line in this semilog-y plot, without any regime switches, since it covers a quiescent period in the signal.
Finally, in panel (i) of Figure 4.2, the spectral radii of the update matrices $\mathbf{D}^{j,n-2}=D^{j,n-2}$ are plotted for $10\leq j\leq n-$ 1, but with the x-axis measured in simulation time units for clarity and illustrative purposes. The y-axis is again on a logarithmic scale. First, it is important to note that for all $j$ and any $n$ (not just the ones shown here), the spectral radii of the update matrices stay below the threshold of $1$, which ensures the convergence of the online smoother as an iterative contraction mapping. Second, the spectral radii, or in this trivial case of one-dimensional latent state space, the absolute value of the update operator $\mathbf{D}^{j,n-2}=D^{j,n-2}$, for every $n$, behaves as an exponentially decreasing function as $j$ decreases down from $n-1$ to 00, which indicates the exponential decrease of the impact region of each new observation by the turbulent dynamics, as discussed in Section 3.2. In contrast, for $n$ values such that $10\leq j\leq n-$ 1 covers an extreme event, the emergence of an instability in the observed signal induces a rapid regime switch and a steep decrease in the spectral radius values, i.e., the influence region of this observation, for when $j$ passes through these intermittent periods, as showed by the steep drops observed at around $t=37.5$ and $t=35$ for each curve covering this period.
Figure 4.2: Panels (a)–(f): Real-time comparison between the filter posterior mean time series (in green) and the one generated from the adaptive-lag online smoother strategy by Theorem 3.1 (in purple). The observed trajectory of $u$ is plotted in panel (a) (in black), while the true trajectory of $v$ is showed in all panels for reference (in blue). Panel (g): Adaptive lag values (measured in time units, i.e., $L_{n}\Delta t$) generated by the algorithm defined in Section 3.4.2, specifically through (3.31)–(3.32). Panel (h): The information gain criterion used to define the adaptive lags through (3.31), for the same time instants as the ones showed in panels (b)–(f), as a function of $j$. These sequences are plotted on a common x-axis for illustration purposes (although their respective $j$’s take different values), for $0\leq j\leq b$, with the x-axis measured in simulation time units. The y-axis is on a logarithmic scale. The threshold value $\delta=10^{-6}$ used in this study is also plotted via a black dashed line for reference. Panel (i): Spectral radii of the update values $D^{j,n-2}$ as a function of $10\leq j\leq n-$ 1 (x-axis is in simulation time units), for the same time instants as the ones displayed in panels (b)–(f). The y-axis is again logarithmically scaled. Panels (h) and (i) share the same legend.
4.2 Lagrangian Data Assimilation of a High-Dimensional Flow Field
Lagrangian data assimilation (LDA) is a specialized method that utilizes the trajectories of moving tracers (e.g., drifters or floaters) as observations used to recover the unknown flow field driving their motion [42, 47, 119, 120, 121, 122]. Unlike Eulerian observations, which are fixed at specific locations, Lagrangian drifters track the movement of fluid parcels over time [123, 124, 125, 126]. This approach is particularly valuable for recovering ocean states in the mid-latitude using floats or in marginal ice zones (MIZ) using sea ice floes as Lagrangian tracers [53, 127, 128]. However, due to the high dimensionality of the state space of the flow field, which is characterized by a large number of spectral modes or high-resolution mesh grids, running a forward-backward offline smoother for state estimation becomes computationally infeasible. Consequently, the adaptive online smoother is essential in practical applications.
The coupled tracer-flow model is given as follows,
${\,\rm d}\mathbf{x}_{\ell}(t)$ $=\mathbf{v}_{\ell}(t){\,\rm d}t+\boldsymbol{\Sigma}^{\mathbf{x}_{\ell}}{\,\rm d}\mathbf{W}_{\mathbf{x}_{\ell}}(t),$ (4.3)
${\,\rm d}\mathbf{v}_{\ell}(t)$ $=\beta\left(\mathbf{u}(t,\mathbf{x}_{\ell})-\mathbf{v}_{\ell}(t)\right){\,\rm d}t+\boldsymbol{\Sigma}^{\mathbf{v}_{\ell}}{\,\rm d}\mathbf{W}_{\mathbf{v}_{\ell}}(t),$ (4.4)
${\,\rm d}\hat{u}_{\mathbf{k}}(t)$ $=(-d_{\mathbf{k}}\hat{u}_{\mathbf{k}}(t)+f_{\mathbf{k}}(t)){\,\rm d}t+\sigma_{\mathbf{k}}{\,\rm d}W_{\mathbf{k}}(t),\qquad\text{with~{}}\mathbf{u}(t,\mathbf{x}_{\ell}(t))=\sum_{\mathbf{k}\in\mathscr{K}}\hat{u}_{\mathbf{k}}(t)e^{i\mathbf{k}\cdot\mathbf{x}_{\ell}(t)}\mathbf{r}_{\mathbf{k}},$ (4.5)
where $\mathbf{x}_{\ell}$ and $\mathbf{v}_{\ell}$ denote the location and velocity of the $\ell$-th tracer, respectively, with $\ell=1,\ldots,L$, while $\mathbf{u}(t,\mathbf{x})$ is the ocean velocity field expressed through a spectral decomposition in its geostrophically balanced or potential vortical incompressible modes, using a Fourier basis in the doubly-periodic spatial domain $[-\pi,\pi]$ [129]. The governing equations of the Fourier coefficients $\hat{u}_{\mathbf{k}}$ are expressed through a set of linear stochastic processes, where $\mathbf{k}\in\mathscr{K}$ denotes the Fourier wavenumber with $\mathscr{K}=[-K,K]^{2}\cap\mathbb{Z}^{2}$ for $K\geq 1$ denoting the two-dimensional discrete lattice collection of Fourier wavenumbers. Note that the stochastic noise mimics the turbulent effects of nonlinearity present in many practical systems, such as the quasi-geostrophic (QG) ocean model. This approximation is justified in many applications [42, 47]. Since the primary goal is to illustrate effectiveness of the online smoother, this simplification is employed in the study here. By treating all the $\mathbf{x}_{\ell}$ as the observational variables and all the $\mathbf{v}_{\ell}$ and $\hat{u}_{\mathbf{k}}$ as the variables for state estimation, the LDA system in (4.3)–(4.5) belongs to the CGNS family. The parameter values in this simulation are the following. The wavenumber bound is $K=2$ such that there are in total 25 modes after excluding the zeroth mode which corresponds to a time-varying random background mean sweep or background velocity field. The total simulation time is $T=5$ with the numerical integration time step (also the observational time frequency) being $\Delta t=0.005$. The total number of tracers is $L=18$. The model parameters are $f_{\mathbf{k}}(t)=0.15e^{\displaystyle 5\pi it}$, $\boldsymbol{\Sigma}^{\mathbf{x}_{\ell}}=0.005\pi\mathbf{I}_{2\times 2}$, $\beta=1$, $\boldsymbol{\Sigma}^{\mathbf{v}_{\ell}}=0.1\mathbf{I}_{2\times 2}$. The damping and noise coefficients $d_{\mathbf{k}}=d_{-\mathbf{k}}$ and $\sigma_{\mathbf{k}}=\sigma_{-\mathbf{k}}$ are randomly drawn from $\mathcal{U}([0.5,1.5])$ and $\mathcal{U}([0.15,0.25])$, respectively, with the conjugacy condition establishing the reality of the underlying velocity field [42, 47].
Figure 4.3 demonstrates the posterior state estimates using the filter (Theorem 2.2) and smoother (Theorem 2.3) in recovering the underlying ocean flow field and the tracer velocity vectors. Note that the amplitudes for the ocean flow field (in blue) and tracer velocity (in the hot colormap) are not shown on the same scale. Panels (a)–-(c) compare the spatial recovery of the ocean state and tracer velocity vectors at $t=2$. Both methods effectively reconstruct the tracer movement in direction and magnitude, with the smoother being slightly more effective. As for the flow velocity field, the smoother exhibits a much higher skill in its recovery compared to the filter solution. This is further illustrated in panels (d)-–(e), where we compare the true time series with the filter and smoother posterior mean states over the observational period for the real part of the ocean Fourier mode $\mathbf{k}=(2,-1)^{\mathtt{T}}$ of the ocean, and the zonal velocity of tracer #1. We can see that the smoother follows true values more closely and with less uncertainty for both cases. This is observed uniformly among all Fourier modes and all tracer velocity vectors.
Figure 4.3: Panel (a): True state of the underlying flow field $t=2$, where the colormap shows the amplitude and the quiver plot represents the velocity field. It is superimposed by the location of the tracers and their velocity vectors. The magnitudes of the quiver plot for the ocean’s velocity field and velocity vectors for the tracers are on distinct scales. Panels (b)–(c): Similar to (a), but corresponding to the estimated state from the posterior mean, calculated through the filter and smoother, respectively. Panels (d)–(e): Comparison between the true time series (in blue) and the posterior mean time series of the filter (in green) and smoother (in red) solutions for the real part of the ocean Fourier mode $\mathbf{k}=(2,-1)^{\mathtt{T}}$ and zonal velocity of tracer #1 (labeled in panels (a)–(c)).
Figure 4.4 compares the computational time (in seconds) and storage requirements (in gigabytes) for the fixed- and adaptive-lag strategies of the online smoother. Additionally, we present the normalized root-mean-square error (NRMSE) of the posterior mean time series to assess the pathwise accuracy of the methods. The normalized RMSE is computed by dividing the RMSE (calculated in the temporal direction) by the standard deviation of the true signal [130, 131, 132], such that the NRMSE indicates an error level comparable to the equilibrium variability. This metric is evaluated as functions of the fixed lag and tolerance parameter $\delta$ and showed on a logarithmically scaled x-axis, where for the latter, a maximum lag bound of $b=300$ is applied throughout and the sequence of entropies is used to define the adaptive lags (i.e., (3.33)). Panels (a) and (b) display fixed-lag results, while panels (d) and (e) display the adaptive-lag strategies. To provide robust results, computational times in panels (b) and (e) are averaged over multiple runs to reduce external fluctuations. For the fixed-lag smoother, both storage and time increase linearly as functions of the fixed-lag, while the NRMSE converges exponentially from the filter solution (zero lag) to the one for the smoother solution (maximum lag), indicating the sufficiency of using an overall short lag to significantly save on computational storage. On the other hand, the storage for the adaptive-lag smoother remains constant with respect to $\delta$ since a fixed upper bound $b$ is being used in all cases, while computational time increases exponentially as $\delta$ approaches zero. The NRMSE for the adaptive-lag method now decreases linearly with tolerance, demonstrating a sharper trade-off between computational resources and error compared to the fixed-lag approach. Panels (c) and (f) further illustrate the recovery skill of these approaches through the temporal average of the information gain for the fixed- and adaptive-lag online smoother strategies beyond the filter solution as functions of the fixed-lag and tolerance parameter, respectively. Both are plotted on an x-axis that is scaled logarithmically and are calculated using the signal-dispersion formula of the relative entropy in (3.26) since both distributions are Gaussian at each time instant. For reference, we show both the signal and dispersion parts, as well as the total information gain. For the fixed-lag smoother, since the zero fixed-lag online smoother coincides with the filter solution, the information gain there is zero, but as the fixed-lag increases, it rapidly and exponentially converges to the information gain value corresponding to the full-backward offline smoother beyond the filter estimates (see Section 3.3). On the other hand, similarly to the results from panels (d) and (e), we have that the temporally-averaged information gain beyond the filter solution is a linear function in $\delta$, where for large values of $\delta$ this is close to zero and linearly approaches the information gain value of the full-backward offline smoother beyond the filter estimates as $\delta$ approaches zero. In both cases, these behaviors are demonstrated in both the signal and dispersion parts and, as such, in the total information gain as well.
Figure 4.4: Panels (a)–(b) & (d)–(e): Analysis of the trade-off between the computational complexity, measured in computational time (in seconds), and required storage capacity (in gigabytes), and lower-order pathwise error statistic of the NRMSE between the posterior mean time series and true signal, for the fixed-lag (panels (a)–(b)) and adaptive-lag online smoother (panels (d)–(e)). All are plotted on a semi-log x-axis, and the x-axis for the adaptive-lag online smoother is reversed for ease of comparison with the corresponding fixed-lag strategy plots. Panels (c) & (f): Temporal average of the information gain for the fixed- (panel (c)) and adaptive-lag (panel (f)) online smoother beyond the filter solution for various values of the fixed lag and tolerance parameter, respectively. Both the signal (in green) and dispersion (in blue) parts of the total information gain (in red) are showed. Their x-axes are logarithmically scaled, and for the adaptive-lag online smoother, it has also reversed direction.
In Figure 4.5, we demonstrate the contrasting behavior of the adaptive-lag online smoother in the LDA problem compared to the dyad-interaction model with intermittent instabilities from Section 4.1. First, in panel (a), similar to panel (h) of Figure 4.2, the information gain criterion used to calculate the adaptive lag values in (3.33) is shown at different time instants for $R_{n}\leq j\leq n-1$ (with the x-axis is measured in simulation time units). In panel (b), same as in panel (i) of Figure 4.2, the spectral radii of the update tensors $\mathbf{D}^{j,n-2}$ are plotted at the same time instants as the sequences in panel (a), for $10\leq j\leq n-$ 1, where for both the y-axis is on a log scale. It is important to observe here the differences and similarities between these plots and the respective ones from Figure 4.2, for the highly intermittent and unstable dyad-interaction model defined in (4.1)–(4.2). In both case studies, these sequences illustrate an exponential behavior in $j$ for all values of $n$. However, for the LDA problem, we can see a uniform rate of increase for these metrics across the board as $j$ increases, due to the parallel lines of constant slope that emerge in both panels. Recall that, in the dyad-interaction numerical experiment, this rate would change depending on whether a rare or extreme event was forming in the observable signal; it could even exhibit regime switches depending on the emergence of intermittent instabilities. This reveals the fact that for the LDA problem of recovering the flow field’s and tracers’ velocities, the observations from the tracers have a uniform impact on the online smoother estimate regardless of when they are observed, which is a result of the stability that the model in (4.3)–(4.5) has, as well as of the fact that the equilibrium distribution of the tracers is nearly uniform due to the incompressibility of the underlying ocean velocity field [42]. This contrasts with the dyad-interaction model, which displays intermittent instabilities due to the latent dynamics acting as anti-damping in the observable ones on rare occasions. Despite this discrepancy, it is noted that the spectral radii of the update tensors $\mathbf{D}^{j,n-2}$, which are high-dimensional matrices in this case, remain below $1$ throughout time and for all values of $n$. This uniform influence region among the observations is also observed in panel (c). Similar to panel (g) in Figure 4.2, it illustrates the adaptive lag values of the online smoother with hyperparameters $b=100$ and $\delta=0.05$ as a function of time, where the adaptive lag is measured in simulation time units (i.e., $L_{n}\Delta t$) and calculated via the sequence of entropies through (3.33). Unlike the adaptive lag values generated from the dyad model in Section 4.1, which highly depend on whether an extreme event is being observed or if we are currently at a period of dormancy, the adaptive lag value here meanders around a constant (about $0.3$ in simulation time units), via uniform, in amplitude, oscillations (after, of course, $n-1\geq b+1$, since for $n-1\leq b$ the adaptive lag value is simply equal to $n-1$ due to (3.32)), and does so steadily in time which indicates the absence of intermittency in the system and the uniform impact each observation has on the online smoother solution.
Figure 4.5: Panel (a): The information gain criterion used to define the adaptive lags through (3.33), as a function of $R_{n}\leq j\leq n-1$, for various values of $n$ (i.e., time instants in $[0,5]$). The x-axis is in simulation time units. The threshold value of interest in this study, $\delta=0.05$, is also plotted via a black dashed line for reference. Panel (b): Spectral radii of the update tensors $D^{j,n-2}$ as a function of $10\leq j\leq n-$ 1 (x-axis is in simulation time units), for the same time instants as the ones displayed in panel (a). Panels (a) and (b) share the same legend. Panel (c): The adaptive lag values (in time units, i.e., $L_{n}\Delta t$) generated for the adaptive-lag online smoother with hyperparameters $b=100$ and $\delta=0.05$ when using (3.33) for their calculation.
4.3 Online Parameter Estimation of Nonlinear Systems with Only Partial Observations
Parameter estimation and model identification for complex nonlinear dynamical systems are crucial for effective state estimation, DA, and forecasting. In partially observed nonlinear systems, parameter and state estimation typically occur simultaneously, for which their solution often requires complicated approximations and expensive numerical methods. Nevertheless, the closed analytic formulas of the online nonlinear smoother (3.11)–(3.12) facilitate an efficient online parameter estimation scheme for the CGNS (2.1)–(2.2) for when only the time series of $\mathbf{x}$ is observed. In this section, we incorporate the adaptive online smoother into an expectation-maximization (EM) algorithm [133, 134, 135, 136, 137] to alternately estimate parameters and hidden states, employing closed analytic formulas for both estimations.
4.3.1 The Online EM Parameter Estimation Algorithm
The basic EM algorithm is outlined in Appendix C, and the mathematical details for the CGNS can be found in [138]. This reference serves as the foundation for developing an online version that utilizes the adaptive-lag online smoother for simultaneous state estimation and model identification, and also includes the tools necessary for incorporating into this algorithm’s skeleton extensions like block decomposition for learning high dimensional systems, sparse identification or regularization, and physical and other constraints on the parameters.
The parameter estimation aims at seeking an optimal estimation of the unknown parameters $\boldsymbol{\theta}$ by maximizing the log-likelihood function corresponding to the partially observed dynamics, also known as the marginal, partial, or conditional log-likelihood. Essentially, our goal is for model identification under the assumption that only a time series from the observable components $\mathbf{x}$ can be obtained, with $\mathbf{y}$ corresponding to the unobserved variables, which nevertheless interacts with $\mathbf{x}$ in (2.1)–(2.2). Since only the time series of $\mathbf{x}$ is observed, the maximum likelihood estimator (MLE) for the marginal log-likelihood is given by
$\hat{\boldsymbol{\theta}}_{\text{MLE}}:=\underset{\boldsymbol{\theta}\in\Theta}{\mathrm{argmax}}\left\{\log\left(p\big{(}\mathbf{x};\boldsymbol{\theta}\big{)}\right)\right\}=\underset{\boldsymbol{\theta}\in\Theta}{\mathrm{argmax}}\left\{\log\left(\int_{\mathbf{y}}p\big{(}\mathbf{x},\mathbf{y};\boldsymbol{\theta}\big{)}{\,\rm d}\mathbf{y}\right)\right\},$ (4.6)
where $\Theta$ denotes the underlying parameter space, which we assume to be a closed and convex subset of $\mathbb{C}^{N}$ for $N\in\mathbb{N}$, and the integration in (4.6), i.e., marginalization, is happening over the state space of $\mathbf{y}$. The EM iteration alternates between performing the so-called expectation step (E-step), which estimates the hidden state of $\mathbf{y}$ using the current estimate for the parameters, and a maximization step (M-step), which computes the parameters that maximize the expected marginal log-likelihood found in the E-step [135, 139, 140]. Since $\mathbf{y}$ is unobserved throughout time, the conditional expectation in the E-step needs to be taken for those measurable functions (measurable with respect to the current observable $\sigma$-algebra) containing $\mathbf{y}$ when computing the log-likelihood estimate, which takes into account the uncertainty in the state estimation through the optimal a-posteriori statistics of $\mathbf{y}$. As is shown in Appendix C, the solution from the E-step is exactly the online smoother posterior distribution.
4.3.2 A Numerical Example
The model we study in this numerical simulation is as follows:
${\,\rm d}u(t)$ $=(-d_{u}u(t)+\gamma u(t)v(t)+F_{u}){\,\rm d}t+\sigma_{u}{\,\rm d}W_{u}(t),$ (4.7)
${\,\rm d}v(t)$ $=(-d_{v}v(t)-\gamma u(t)^{2}+F_{v}){\,\rm d}t+\sigma_{v}{\,\rm d}W_{v}(t),$ (4.8)
where $d_{u},d_{v},\gamma_{1},\gamma_{2},F_{u},F_{v}\in\mathbb{R}$ and $\sigma_{u},\sigma_{v}>0$, with $W_{u}$ and $W_{v}$ being two mutually independent Wiener processes. The system (4.7)–(4.8) fits the CGNS framework for $\mathbf{x}:=u$ and $\mathbf{y}=v$. Similar to the model in (4.1)–(4.2) of Section 4.1, the variable $v$ acts as a stochastic damping in the observable process, which leads to intermittent extreme events. The effect of these instabilities in the convergence of the algorithm is especially studied. We assume a time series of $u$ is continuously observed while there is no observation of $v$. To simplify the study in this section, we assume that the noise feedbacks $\sigma_{u}$ and $\sigma_{v}$ are known, although our framework does not require such a restriction [138]. Thus, the parameter vector for this model is as follows:
$\boldsymbol{\theta}:=(d_{u},\gamma_{1},F_{u},d_{v},\gamma_{2},F_{v})^{\mathtt{T}}.$ (4.9)
The true parameter values are the following:
$d_{u}=1,\quad\gamma=3,\quad F_{u}=1,\quad\sigma_{u}=0.5,$
$d_{v}=1,\quad F_{v}=0.8,\quad\sigma_{v}=1.$
The numerical setup is as follows. The initial guess of the parameter values is $\boldsymbol{\theta}_{0}=(2,6,2,0.5,6,0.4)^{\mathtt{T}}$. The numerical integration time step, which is also the observational frequency in such a continuous observational case, is $\Delta t=0.001$. The total length of observations is $T=200$ time units. Since implementing the online smoother requires a certain history of previous state estimation solutions, the observed time series of $u$ within a short initial window $[0,T_{\text{ini}}]$, with $T_{\text{ini}}=10$, is utilized to initially carry out the filtering solution and online smoother estimates, based on $\boldsymbol{\theta}_{0}$. This can be treated as a burn-in or learning period for the algorithm, and also aids in avoiding the possibility of short-time blowup. Afterwards, with the arrival of each new observation, we update the adaptive-lag online smoother state estimations via the procedure in (3.27)–(3.28) and then re-estimate the parameters using the MLE. The update happens at every observation as to ensure stability. The hyperparameters in the adaptive-lag method are $b=1000$ (where $b\Delta t=1$ time unit) and $\delta=10^{-4}$, where we use the actual sequence of information gains to define the adaptive lag at each new observation, i.e., (3.33).
Figure 4.6 presents the results from using the online EM algorithm. Panels (a) and (b) display the true trajectories of the observed and unobserved variables, $u$ and $v$, respectively, displayed in blue. These trajectories are generated by running equations (4.7)–(4.8) forward in time. Alongside the true trajectories, the time series obtained by simulating the model with the estimated parameters at the final iteration of the online EM algorithm are included in red. Note that the same values for the Wiener processes are used for these two signals, allowing us to have a point-wise comparison. The recovery of the signal for both observable and unobserved variables is notably skillful. It is worth highlighting that, in panel (b), the dashed horizontal lines indicate the true (in black) and online EM estimated (in green) anti-damping threshold lines, above which the unobserved variable effectively acts as anti-damping in the observable process, initiating intermittent instabilities in the signal. Their consistency is critical for retrieving the true dynamics using the estimated parameters. Panel (c) shows the adaptive lag values, measured in time units (i.e., $L_{n}\Delta t$) for the online smoother, illustrating that longer lags are typically required during the onset of the intermittent phases as to effectively capture the dynamics. The efficacious recovery of the true state is also demonstrated through higher-order statistics, such as the PDFs and ACFs for both $u$ (in panels (d)–(e)) and $v$ (in panels (f)–(g)). Finally, panels (h)–(m) present the trace plots produced by the online EM iterations for various parameters of interest. The parameters in the observed process $u$ are nearly perfectly recovered, including the nonlinearity feedback $\gamma$ and the deterministic constant forcing $F_{v}$ in the process of the unobserved variable. The only parameter that is not perfectly recovered is the damping $d_{u}$ in the process of the unobserved variable, with the algorithm converging to twice its true value. This is expected, as these parameters, along with the noise feedback, are typically the hardest to estimate due to observability issues, where their contribution to the observed process is relatively weak. Nevertheless, this inaccuracy does not hinder the model with the estimated parameters from reproducing key dynamical and statistical features, as shown in panels (a)–(b) and (d)–(g).
An important observation from the trace plots is that extreme events in the observable signal enhance the convergence skill of the online EM algorithm within the CGNS framework. In panel (a), four extreme events are marked as A, B, C, and D, characterized by a high signal-to-noise ratio. The trace plots in panels (h)–(m) demonstrate that extreme events influence parameter convergence. The high signal-to-noise ratio mitigates the uncertainty in observations and results in strong observability of the system, which naturally facilitates parameter estimation [141]. This is particularly evident in the anti-damping threshold (panel (m)), highlighting the role of these events in identifying the intermittent nature of the underlying dynamics.
Figure 4.6: Panel (a): True trajectory of the observable variable $u$ over the simulation period (in blue), and the recovered trajectory from using the same white noise values used to generate the true signal (by utilizing the same seed number) but where the model in (4.7)–(4.8) is instead defined through the recovered parameter values at the last iteration of the online EM algorithm (in red). The dashed rectangular box denotes the period of running the filter and online smoother algorithms with $\boldsymbol{\theta}_{0}$, which is $[0,10]$. Four extreme events are marked in the observed time series (denoted by A, B, C, and D). Panel (b): Same as (a) but for the unobserved variable $v$. The true and online EM estimated anti-damping threshold $d_{u}/\gamma_{1}$ are marked by dashed lines, in black and green, respectively. Panel (c): Adaptive lag values generated using (3.33) (measured in time units, i.e., $L_{n}\Delta t$), as a function of simulation time. Panels (d)–(g): PDFs of $u$ (panel (d)) and $v$ (panel (f)) and their ACFs (panels (e) and (g), respectively), for both the true model parameters and the ones obtained at the last iteration of the online EM algorithm. Panels (h)–(m): Trace plots produced by the online EM algorithm.
5 Conclusions
In this paper, a computationally efficient algorithm for online adaptive-lag optimal smoother-based state estimation with partial observations is developed. Closed analytical formulae are available for this online smoother. It applies to CGNS, representing a rich class of CTNDSs with wide applications in neuroscience, ecology, atmospheric science, geophysics, and many other fields. The adaptive-lag strategy allows a significant reduction of computational storage. Notably, an information-theoretic criterion, exploiting uncertainty reduction across the entire posterior distribution, is developed to calculate the adaptive lag value, which helps in preserving dynamical and statistical consistency.
The adaptive-lag online smoother has been applied to systems with intermittent instabilities (Section 4.1). It effectively recovers extreme events and discovers the causal dependence between different state variables. It is also applicable to high-dimensional systems, such as the LDA problem. The state estimation based on such an online smoother shows nearly the same accuracy as the standard forward-backward smoother while requiring significantly less computational storage (Section 4.2). In addition, the online adaptive-lag smoother facilitates real-time parameter estimation. The study highlights the importance of utilizing observed extreme events to accelerate the convergence of parameter estimation (Section 4.3).
Potential future research topics in this direction are as follows. The closed-form expressions of the online smoother estimates can aid in assessing extreme events and their impact on the CGNS dynamics, as well as in identifying the most sensitive directions that exhibit rapid fluctuations during specific events using information-theoretic methods [98]. This is facilitated by the fact that the CGNS framework enjoys analytic formulae for such conditional distributions [36]. Second, the adaptive-lag online smoother can serve as a useful tool for causal inference and information flow analysis. On the one hand, one can exploit intermittent extreme events to improve state estimation, which helps reduce the model error in developing parsimonious surrogate models via causal inference. On the other hand, the conditional Gaussianity of the framework allows the application of transfer or causation entropy to explore these topics in future studies. Finally, future work also involves developing a rigorous investigation of the spectral properties of the update tensors $\mathbf{D}^{j,n-2}$ used in the online smoother, as well as of the matrices $\mathbf{E}^{j}$ that define them, potentially through suitable concentration inequalities. This includes studying the temporal behavior of the information-theoretic criterion for adaptive lags, $\mathcal{P}\left(p^{j,n}_{\text{updated}},p^{j,n}_{\text{lagged}}\right)$ (and its LSDF). Numerical case studies have shown that these quantities increase exponentially over time, with the rate depending on whether a quiescent or intermittent signal is observed. Notably, the spectrum of $\mathbf{D}^{j,n-2}$ remains below the convergence threshold of $1$, which is reassuring. A rigorous analysis of these properties is a natural future topic.
Acknowledgments
N.C. is grateful to acknowledge the support of the Office of Naval Research (ONR) N00014-24-1-2244 and Army Research Office (ARO) W911NF-23-1-0118. M.A. is supported as a research assistant under these grants.
Appendix A Sufficient Assumptions for the Results of the CGNS Framework
Let the following denote the standard elements of the vector- and matrix-valued functionals appearing in (2.1)–(2.2) as model parameters:
$\mathbf{f}^{\mathbf{x}}(t,\mathbf{x}):=(f^{\mathbf{x}}_{1}(t,x),\ldots,f^{\mathbf{x}}_{k}(t,x))^{\mathtt{T}},\quad\mathbf{f}^{\mathbf{y}}(t,\mathbf{x}):=(f^{\mathbf{y}}_{1}(t,x),\ldots,f^{\mathbf{y}}_{l}(t,x))^{\mathtt{T}},$
$\boldsymbol{\Lambda}^{\mathbf{x}}(t,\mathbf{x}):=\left(\Lambda^{\mathbf{x}}_{ij}(t,\mathbf{x})\right)_{k\times l},\quad\boldsymbol{\Sigma}_{1}^{\mathbf{x}}(t,\mathbf{x}):=\left(\Sigma^{x,1}_{ij}(t,\mathbf{x})\right)_{k\times d},\quad\boldsymbol{\Sigma}_{2}^{\mathbf{x}}(t,\mathbf{x}):=\left(\Sigma^{x,2}_{ij}(t,\mathbf{x})\right)_{k\times r},$
$\boldsymbol{\Lambda}^{\mathbf{y}}(t,\mathbf{x}):=\left(\Lambda^{\mathbf{y}}_{ij}(t,\mathbf{x})\right)_{l\times l},\quad\boldsymbol{\Sigma}_{1}^{\mathbf{y}}(t,\mathbf{x}):=\left(\Sigma^{y,1}_{ij}(t,\mathbf{x})\right)_{l\times d},\quad\boldsymbol{\Sigma}_{2}^{\mathbf{y}}(t,\mathbf{x}):=\left(\Sigma^{y,2}_{ij}(t,\mathbf{x})\right)_{l\times r}.$
As to be able to obtain the main results which define the CGNS framework and its potency in DA through closed-form expressions for the posterior statistics, a set of sufficient regularity conditions needs to be assumed a-priori. For the following conditions, each respective pair of indices $i$ and $j$ takes all admissible values and $\mathbf{z}$ is a $k$-dimensional function in $C^{0}([0,T];\mathbb{C}^{k})$:
  1. (1)
    We assume:
    $\int_{0}^{T}\Big{[}$ $|f^{\mathbf{y}}_{i}(t,\mathbf{z})|+|\Lambda^{\mathbf{y}}_{ij}(t,\mathbf{z})|+\big{|}f^{\mathbf{x}}_{i}(t,\mathbf{z})\big{|}^{2}+\big{|}\Lambda^{\mathbf{x}}_{ij}(t,\mathbf{z})\big{|}^{2}$
    $+\big{|}\Sigma^{x,1}_{ij}(t,\mathbf{z})\big{|}^{2}+\big{|}\Sigma^{x,2}_{ij}(t,\mathbf{z})\big{|}^{2}+\big{|}\Sigma^{y,1}_{ij}(t,\mathbf{z})\big{|}^{2}+\big{|}\Sigma^{y,2}_{ij}(t,\mathbf{z})\big{|}^{2}\Big{]}\mathrm{d}t<+\infty.$
    This ensures the existence of the integrals in (2.1)–(2.2) [35, 36].
  2. (2)
    If $g(t,\mathbf{z})$ denotes any element of the multiplicative factors in the drift dynamics, $\boldsymbol{\Lambda}^{\mathbf{x}}$ and $\boldsymbol{\Lambda}^{\mathbf{y}}$, or of the noise feedback matrices, $\boldsymbol{\Sigma}_{m}^{\mathbf{x}}(t,\mathbf{z})$ and $\boldsymbol{\Sigma}_{m}^{\mathbf{y}}(t,\mathbf{z})$ for $m=1,2$, and $K(s)$ is a nondecreasing right-continuous function taking values in $[0,1]$, then there $\exists L_{2},L_{3},L_{4},L_{5}>0$ such that for any $\mathbf{z}$ and $\mathbf{w}$ $k$-dimensional functions in $C^{0}([0,T];\mathbb{C}^{k})$ we have
    $|g(t,\mathbf{z})-g(t,\mathbf{w})|^{2}\leq L_{2}\int_{0}^{t}\left\|\mathbf{z}(s)-\mathbf{w}(s)\right\|_{2}^{2}\mathrm{d}K(s)+L_{3}\left\|\mathbf{z}(t)-\mathbf{w}(t)\right\|_{2}^{2},\ \forall t\in[0,T],$
    $|g(t,\mathbf{z})|^{2}\leq L_{4}\int_{0}^{t}(1+\left\|\mathbf{z}(s)\right\|_{2}^{2})\mathrm{d}K(s)+L_{5}(1+\left\|\mathbf{z}(t)\right\|_{2}^{2}),\ \forall t\in[0,T].$
    These integrals are to be understood in the Lebesgue–Stieltjes integration sense.
  3. (3)
    The sum of the Gramians (with respect to rows) of the noise coefficient matrices in the observable process are uniformly nonsingular, i.e., the elements of the inverse of
    $(\boldsymbol{\Sigma}^{\mathbf{x}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})(t,\mathbf{z})=\boldsymbol{\Sigma}_{1}^{\mathbf{x}}(t,\mathbf{z})\boldsymbol{\Sigma}_{1}^{\mathbf{x}}(t,\mathbf{z})^{\dagger}+\boldsymbol{\Sigma}_{2}^{\mathbf{x}}(t,\mathbf{z})\boldsymbol{\Sigma}_{2}^{\mathbf{x}}(t,\mathbf{z})^{\dagger},$
    are uniformly bounded in $[0,T]$.
  4. (4)
    $\int_{0}^{T}\mathbb{E}\left[|\Lambda^{\mathbf{x}}_{ij}(t,\mathbf{z}(t))y_{j}(t)|\right]\mathrm{d}t<+\infty$.
  5. (5)
    $\mathbb{E}\left[|y_{j}(t)|\right]<+\infty,\ t\in[0,T]$.
  6. (6)
    For ${\mu_{\text{f}}}_{j}(t):=\mathbb{E}\left[y_{j}(t)\big{|}\mathcal{F}_{t}^{\mathbf{x}}\right]$, where $t\in[0,T]$ and $j=1,\ldots,l$, we assume that
    $\mathbb{P}\left(\int_{0}^{T}\left|\Lambda^{\mathbf{x}}_{ij}(t,\mathbf{z}(t))\mu_{\text{f},j}(t)\right|^{2}\mathrm{d}t<+\infty\right)=1.$
Accompanying the conditional Gaussianity provided by Theorem 2.1, we further assume the following additional conditions (which concur with assumptions (1)(6) from before):
  • (7)
    $|\Lambda^{\mathbf{x}}_{ij}(t,\mathbf{z})|,|\Lambda^{\mathbf{y}}_{ij}(t,\mathbf{z})|\leq L_{1}$ for some (uniform) $L_{1}>0$, $\forall t\in[0,T]$.
  • (8)
    $\int_{0}^{T}\mathbb{E}\left[|\boldsymbol{\Lambda}_{ij}^{\mathbf{x}}(t,\mathbf{x})|^{4}+|\mathbf{f}_{i}^{\mathbf{y}}(t,\mathbf{x})|^{4}+|\Sigma^{y,1}_{ij}(t,\mathbf{x})|^{4}+|\Sigma^{y,2}_{ij}(t,\mathbf{x})|^{4}\right]\mathrm{d}t<+\infty$.
  • (9)
    $\mathbb{E}\left[\left\|\mathbf{y}(0)\right\|_{2}^{4}\right]<+\infty$.
Finally, as to be able to derive discrete-time smoother equations which are consistent with their continuous-time counterparts in the limit as their associated temporal partition’s mesh size converges to zero, with this recovery being first order-accurate under an explicit Euler-Maruyama numerical scheme, we assume that:
  • (10)
    $\mathbb{E}\left[\|\mathbf{x}(0)\|_{2}^{2}+\|\mathbf{y}(0)\|_{2}^{2}\right]<+\infty$. This also implies $\mathbb{P}$-a.s. finiteness of the initial conditions by Markov’s inequality.
  • (11)
    For $\mathbf{z}\in C^{0}([0,T];\mathbb{C}^{k})$, $\mathbf{u}\in C^{0}([0,T];\mathbb{C}^{l})$, and $\mathbf{f}:=\begin{pmatrix}\mathbf{0}_{k\times k}&\boldsymbol{\Lambda}^{\mathbf{x}}\\ \mathbf{0}_{l\times k}&\boldsymbol{\Lambda}^{\mathbf{y}}\end{pmatrix}\begin{pmatrix}\mathbf{x}\\ \mathbf{y}\end{pmatrix}+\begin{pmatrix}\mathbf{f}^{\mathbf{x}}\\ \mathbf{f}^{\mathbf{y}}\end{pmatrix}$, $\boldsymbol{\Sigma}:=\begin{pmatrix}\boldsymbol{\Sigma}_{1}^{\mathbf{x}}&\boldsymbol{\Sigma}_{2}^{\mathbf{x}}\\ \boldsymbol{\Sigma}_{2}^{\mathbf{y}}&\boldsymbol{\Sigma}_{2}^{\mathbf{y}}\end{pmatrix}$, there exist $L_{6},L_{7}>0$ such that
    $\left\|\mathbf{f}(t,\mathbf{z},\mathbf{u})-\mathbf{f}(s,\mathbf{z},\mathbf{u})\right\|_{2}+\left\|\boldsymbol{\Sigma}(t,\mathbf{z})-\boldsymbol{\Sigma}(s,\mathbf{z})\right\|_{2}\leq L_{6}\big{(}1+(\left\|\mathbf{z}\right\|^{2}_{2}+\left\|\mathbf{u}\right\|^{2}_{2})^{1/2}\big{)}|t-s|^{1/2},\ \forall t,s\in[0,T].$
[…]
      We implicitly assume that condition (10) from Andreou & Chen [68] is satisfied for any stochastic discrete-time approximation which recovers the strong solution to the CGNS at vanishing mesh size, and specifically for the numerical solution provided by the Euler-Maruyama scheme. In other words, we assume that the initial distributions are not being severely misidentified.
  • (12)
    Adopting a uniform temporal discretization of $[0,T]$, with $\Delta t=T/J$ being its norm or mesh and $t_{j}=j\Delta t$ for $j=0,1,\ldots,J$ being its nodes, where $J\in\mathbb{N}$, we assume that
    $a_{\Delta t}(M):=\underset{\theta>0}{\inf}\Bigg{\{}\mathbb{E}\Bigg{[}e^{\theta\big{(}M\big{(}1+\underset{0\leq j\leq J}{\max}\big{\{}\big{\|}\boldsymbol{\Sigma}_{1}^{\mathbf{y},j}\big{\|}_{2}^{2}+\big{\|}\boldsymbol{\Sigma}_{2}^{\mathbf{y},j}\big{\|}_{2}^{2}\big{\}}\big{)}-1/\Delta t\big{)}}\Bigg{]}\Bigg{\}}=o(\Delta t),$
    as $\Delta t\to 0^{+}$, for every $M=M(\omega)$ with $\mathbb{P}(M<+\infty)=1$.
The optimal discrete smoother equations derived under these additional sufficient assumptions are essential in establishing a corresponding optimal online discrete smoother which runs forward in time.
Appendix B Proof of the Optimal Online Forward-In-Time Discrete Smoother
In this appendix we outline the details of the proof to Theorem 3.1 which outlines the procedure for the online forward-in-time discrete smoother for optimal state estimation of the unobserved processes.
Proof of Theorem 3.1.
From the derivations in Andreou & Chen [68] concerning the discrete smoother, for $n\in\mathbb{N}$ and $j=0,1,\ldots,n-1$, we have that
$\boldsymbol{\mu}_{\text{{s}}}^{j,n}$ $=\boldsymbol{\mu}_{\text{{f}}}^{j}+\mathbf{E}^{j}\left(\boldsymbol{\mu}_{\text{{s}}}^{j+1,n}-(\mathbf{I}_{l\times l}+\boldsymbol{\Lambda}^{\mathbf{y},j}\Delta t)\boldsymbol{\mu}_{\text{{f}}}^{j}-\mathbf{f}^{\mathbf{y},j}\Delta t\right)+\mathbf{F}^{j}\left(\mathbf{x}^{j+1}-\mathbf{x}^{j}-(\boldsymbol{\Lambda}^{\mathbf{x},j}\boldsymbol{\mu}_{\text{{f}}}^{j}+\mathbf{f}^{\mathbf{x},j})\Delta t\right)=\mathbf{E}^{j}\boldsymbol{\mu}_{\text{{s}}}^{j+1,n}+\mathbf{b}^{j},$ (B.1)
$\mathbf{R}_{\text{{s}}}^{j,n}$ $=\mathbf{R}_{\text{{f}}}^{j}-\mathbf{C}_{j+1}^{j}\mathbf{C}^{j}_{22}(\mathbf{C}_{j+1}^{j})^{\dagger}+\mathbf{E}^{j}\mathbf{R}_{\text{{s}}}^{j+1,n}(\mathbf{E}^{j})^{\dagger}=\mathbf{E}^{j}\mathbf{R}_{\text{{s}}}^{j+1,n}(\mathbf{E}^{j})^{\dagger}+\mathbf{P}_{j+1}^{j},$ (B.2)
where
$\mathbf{b}^{j}$ $:=\boldsymbol{\mu}_{\text{f}}^{j}-\mathbf{E}^{j}\left((\mathbf{I}_{l\times l}+\boldsymbol{\Lambda}^{\mathbf{y},j}\Delta t)\boldsymbol{\mu}_{\text{{f}}}^{j}+\mathbf{f}^{\mathbf{y},j}\Delta t\right)+\mathbf{F}^{j}\left(\mathbf{x}^{j+1}-\mathbf{x}^{j}-(\boldsymbol{\Lambda}^{\mathbf{x},j}\boldsymbol{\mu}_{\text{{f}}}^{j}+\mathbf{f}^{\mathbf{x},j})\Delta t\right),$
$\mathbf{P}^{j}_{j+1}$ $:=\mathbf{R}_{\text{f}}^{j}-\mathbf{E}^{j}(\mathbf{I}_{l\times l}+\boldsymbol{\Lambda}^{\mathbf{y},j}\Delta t)\mathbf{R}_{\text{f}}^{j}-\mathbf{F}^{j}\boldsymbol{\Lambda}^{\mathbf{x},j}\mathbf{R}_{\text{f}}^{j}\Delta t,$
which immediately shows that (3.14) and (3.15) hold by plugging-in $j=n-1$ into these, since the smoother and filter estimates are the same at the end point, i.e., $\boldsymbol{\mu}_{\text{s}}^{n,n}=\boldsymbol{\mu}_{\text{f}}^{n}$ and $\mathbf{R}_{\text{s}}^{n,n}=\mathbf{R}_{\text{f}}^{n}$. As such, we just need to prove (3.11) and (3.12).
We start by proving the following relation for $n\in\mathbb{N}$ and $j=0,1,\ldots,n-1$:
$\mathbf{D}^{j,n-1}=\mathbf{E}^{j}\mathbf{E}^{j+1}\cdots\mathbf{E}^{n-1}=:\overset{{\curvearrowright}}{\prod^{n-1}_{i=j}}\mathbf{E}^{i}=\mathbf{E}^{j}\mathbf{D}^{j+1,n-1}.$ (B.3)
The curved arrow pointing to the right above the product symbol indicates the order with which we expand said product since we are dealing with matrices, thus making the product, in general, noncommutative. Also, note that (B.3) is able to encompass both the second and third relations of (3.13), but not the first one due to its irregular form. Nonetheless, it could be extraneously incorporated in (B.3) by noting that if $j=n$, and so the product on the right-hand side is “empty”, then $\mathbf{D}^{n,n-1}$ is trivially defined as the identity matrix $\mathbf{I}_{l\times l}$.
For $n\in\mathbb{N}_{\geq 2}$ and $j=0,1,\ldots,n-1$, we have from iterative application of the appropriate relation in (3.13) (the second and third relations are appropriately applied in what follows, possibly after some reindexing, without loss of generality), that
$\mathbf{D}^{j,n-1}=\mathbf{D}^{j,n-2}\mathbf{E}^{n-1}$ =$=\begin{cases}\mathbf{E}^{n-1},&\text{ for }j=n-1\\ \mathbf{E}^{n-2}\mathbf{E}^{n-1},&\text{ for }j=n-2\\ \mathbf{D}^{j,n-3}\mathbf{E}^{n-2}\mathbf{E}^{n-1},&\text{ for }j<n-2=\begin{cases}\mathbf{E}^{n-3}\mathbf{E}^{n-2}\mathbf{E}^{n-1},&\text{ for }j=n-3\\ \mathbf{D}^{j,n-4}\mathbf{E}^{n-3}\mathbf{E}^{n-2}\mathbf{E}^{n-1},&\text{ for}j<n-3\end{cases}\end{cases}$
$=\cdots=\mathbf{D}^{j,j}\mathbf{E}^{j+1}\cdots\mathbf{E}^{n-3}\mathbf{E}^{n-2}\mathbf{E}^{n-1}\equiv\overset{{\curvearrowright}}{\prod^{n-1}_{i=j}}\mathbf{E}^{i},$
which proves the relation on the left-hand side of (B.3). Similarly, from the previous result and (3.13) (the first relation), we have
$\mathbf{E}^{j}\mathbf{D}^{j+1,n-1}=\begin{cases}\mathbf{E}^{j},&\text{ for }j+1=n\\ \mathbf{E}^{j}\overset{{\curvearrowright}}{\prod^{n-1}_{i=j+1}}\mathbf{E}^{i},&\text{ for }j+1\leq n-1<n\end{cases}\equiv\overset{{\curvearrowright}}{\prod^{n-1}_{i=j}}\mathbf{E}^{i},$
which is the same as the right-hand side of (B.3). For $n=1$, and so $j=0$ necessarily, we have
$\mathbf{D}^{0,0}=\mathbf{E}^{0}=\overset{{\curvearrowright}}{\prod^{0}_{i=0}}\mathbf{E}^{i},$
by the second relation in (3.13), while by the first relation in (3.13) we have
$\mathbf{E}^{0}\mathbf{D}^{1,0}=\mathbf{E}^{0}=\overset{{\curvearrowright}}{\prod^{0}_{i=0}}\mathbf{E}^{i},$
and so by combining these two results proves (B.3) also for $n=1$ and $j=0$. Collecting all these results together, establishes (B.3) for $n\in\mathbb{N}$ and $j=0,\ldots,n-1$.
We proceed now by proving that for $n\in\mathbb{N}$ and $10\leq j\leq n-$ 1, the following equations are valid:
$\boldsymbol{\mu}_{\text{s}}^{j,n}$ $=\mathbf{D}^{j,n-1}\boldsymbol{\mu}_{\text{f}}^{n}+\mathbf{b}^{j}+\sum_{r=j+1}^{n-1}\mathbf{D}^{j,r-1}\mathbf{b}^{r},$ (B.4)
$\mathbf{R}_{\text{s}}^{j,n}$ $=\mathbf{D}^{j,n-1}\mathbf{R}_{\text{f}}^{n}(\mathbf{D}^{j,n-1})^{\dagger}+\mathbf{P}_{j+1}^{j}+\sum_{r=j+1}^{n-1}\mathbf{D}^{j,r-1}\mathbf{P}_{r+1}^{r}(\mathbf{D}^{j,r-1})^{\dagger}.$ (B.5)
We start with the equation for the smoother mean. We have by an iterative application of (B.1) for $n\in\mathbb{N}$ and $j=0,\ldots,n-1$ that
$\boldsymbol{\mu}_{\text{s}}^{j,n}$ $=\mathbf{E}^{j}\underbrace{\boldsymbol{\mu}_{\text{s}}^{j+1,n}}_{\displaystyle\mathclap{\hskip 253.22934pt=\mathbf{E}^{j+1}\boldsymbol{\mu}_{\text{s}}^{j+2,n}+\mathbf{b}^{j+1},\text{ if }j+1<n,\text{ otherwise equals }\boldsymbol{\mu}_{\text{f}}^{n}}}+\mathbf{b}^{j}$
[…]
$=\cdots\ \left(\text{Likewise where at the }k\text{-th step we have }\boldsymbol{\mu}_{\text{s}}^{j+k,n}=\begin{cases}\mathbf{E}^{j+k}\boldsymbol{\mu}_{\text{s}}^{j+k+1,n}+\mathbf{b}^{j+k},&\text{ for }j+k<n\\ \boldsymbol{\mu}_{\text{f}}^{n,n},&\text{ for }j+k=n\end{cases}\right)$
[…]
$=\mathbf{E}^{j}\mathbf{E}^{j+1}\cdots\mathbf{E}^{n-1}\boldsymbol{\mu}_{\text{f}}^{n}+\mathbf{b}^{j}+\left(\mathbf{E}^{j}\mathbf{b}^{j+1}+\mathbf{E}^{j}\mathbf{E}^{j+1}\mathbf{b}^{j+2}+\cdots+\left(\overset{{\curvearrowright}}{\prod^{n-2}_{i=j}}\mathbf{E}^{i}\right)\mathbf{b}^{n-1}\right).$
Using now (B.3) (possibly after some reindexing, without loss of generality), then this expression reduces down to
$\boldsymbol{\mu}_{\text{s}}^{j,n}=\mathbf{D}^{j,n-1}\boldsymbol{\mu}_{\text{f}}^{n}+\mathbf{b}^{j}+\sum_{r=j+1}^{n-1}\mathbf{D}^{j,r-1}\mathbf{b}^{r},$
which is exactly what we wanted to prove for the smoother mean in (B.4). We now turn our attention to the smoother covariance matrix. In a very similar fashion, now by iteratively using (B.2) for $n\in\mathbb{N}$ and $j=0,\ldots,n-1$, we get
$\mathbf{R}_{\text{s}}^{j,n}$ $=\mathbf{E}^{j}\underbrace{\mathbf{R}_{\text{s}}^{j+1,n}}_{\displaystyle\mathclap{\hskip 295.90848pt=\mathbf{E}^{j+1}\mathbf{R}_{\text{s}}^{j+2,n}(\mathbf{E}^{j+1})^{\dagger}+\mathbf{P}^{j+1}_{j+2},\text{ if }j+1<n,\text{ otherwise equals }\mathbf{R}_{\text{f}}^{n}}}(\mathbf{E}^{j})^{\dagger}+\mathbf{P}^{j}_{j+1}$
[…]
$=\cdots\ \left(\text{Likewise where at the }k\text{-th step we have }\boldsymbol{\mu}_{\text{s}}^{j+k,n}=\begin{cases}\mathbf{E}^{j+k}\boldsymbol{\mu}_{\text{s}}^{j+k+1,n}+\mathbf{b}^{j+k},&\text{ for }j+k<n\\ \boldsymbol{\mu}_{\text{f}}^{n,n},&\text{ for }j+k=n\end{cases}\right)$
[…]
$=\mathbf{E}^{j}\mathbf{E}^{j+1}\cdots\mathbf{E}^{n-1}\mathbf{R}_{\text{f}}^{n,n}(\mathbf{E}^{n-1})^{\dagger}\cdots(\mathbf{E}^{j+1})^{\dagger}(\mathbf{E}^{j})^{\dagger}+\mathbf{P}_{j+1}^{j}$
$\hskip 56.9055pt+\left(\mathbf{E}^{j}\mathbf{P}^{j+1}_{j+2}(\mathbf{E}^{j})^{\dagger}+\mathbf{E}^{j}\mathbf{E}^{j+1}\mathbf{P}^{j+2}_{j+3}(\mathbf{E}^{j+1})^{\dagger}(\mathbf{E}^{j})^{\dagger}+\cdots+\left(\overset{{\curvearrowright}}{\prod^{n-2}_{i=j}}\mathbf{E}^{i}\right)\mathbf{P}^{n-1}_{n}\left(\overset{{\curvearrowleft}}{\prod^{n-2}_{i=j}}(\mathbf{E}^{i})^{\dagger}\right)\right).$
Like we did before, by using (B.3) (possibly after some reindexing, without loss of generality), and by noticing that
$(\mathbf{D}^{j,r-1})^{\dagger}=\left(\overset{{\curvearrowright}}{\prod^{r-1}_{i=j}}\mathbf{E}^{i}\right)^{\dagger}=\overset{{\curvearrowleft}}{\prod^{r-1}_{i=j}}(\mathbf{E}^{i})^{\dagger},$
for $r=n,n-1,\ldots,j+1$, then the previous expression is equivalent to,
$\mathbf{R}_{\text{s}}^{j,n}=\mathbf{D}^{j,n-1}\mathbf{R}_{\text{f}}^{n}(\mathbf{D}^{j,n-1})^{\dagger}+\mathbf{P}_{j+1}^{j}+\sum_{r=j+1}^{n-1}\mathbf{D}^{j,r-1}\mathbf{P}_{r+1}^{r}(\mathbf{D}^{j,r-1})^{\dagger},$
which is exactly what we wanted to prove for the smoother covariance in (B.5).
Now to verify the recursive formulae given in (3.11) and (3.12), we use (B.4) and (B.5) by looking at the differences $\boldsymbol{\mu}_{\text{s}}^{j,n}-\boldsymbol{\mu}_{\text{s}}^{j,n-1}$ and $\mathbf{R}_{\text{s}}^{j,n}-\mathbf{R}_{\text{s}}^{j,n-1}$, respectively. First, observe that (3.11) and (3.12) trivially hold in the case where $n=1$, and as such we necessarily have $j=0,\ldots,n-1\Leftrightarrow j=0$. This is immediate by the first relation in (3.13) and the fact that $\boldsymbol{\mu}_{\text{s}}^{0,0}=\boldsymbol{\mu}_{\text{f}}^{0}$ and $\mathbf{R}_{\text{s}}^{0,0}=\mathbf{R}_{\text{f}}^{0}$. As such, we only consider the case where $n\in\mathbb{N}_{\geq 2}$ and $j=0,\ldots,n-1$. From (B.4) we have,
$\boldsymbol{\mu}_{\text{s}}^{j,n}-\boldsymbol{\mu}_{\text{s}}^{j,n-1}=\mathbf{D}^{j,n-1}\boldsymbol{\mu}_{\text{f}}^{n}+\mathbf{D}^{j,n-2}\mathbf{b}^{n-1}-\mathbf{D}^{j,n-2}\boldsymbol{\mu}_{\text{f}}^{n-1},$
and by (B.5) we get,
$\mathbf{R}_{\text{s}}^{j,n}-\mathbf{R}_{\text{s}}^{j,n-1}=\mathbf{D}^{j,n-1}\mathbf{R}_{\text{f}}^{n}(\mathbf{D}^{j,n-1})^{\dagger}+\mathbf{D}^{j,n-2}\mathbf{P}_{n}^{n-1}(\mathbf{D}^{j,n-2})^{\dagger}-\mathbf{D}^{j,n-2}\mathbf{R}_{\text{f}}^{n-1}(\mathbf{D}^{j,n-2})^{\dagger}.$
Replacing now the coefficient $\mathbf{D}^{j,n-1}$ by $\mathbf{D}^{j,n-2}\mathbf{E}^{n-1}$ via the third relation in (3.13), since it holds for $n\in\mathbb{N}_{\geq 2}$, $j=0,\ldots,n-1$, for the smoother mean we have,
[…]
$=\boldsymbol{\mu}_{\text{s}}^{j,n-1}+\mathbf{D}^{j,n-2}\left(\boldsymbol{\mu}_{\text{s}}^{n-1,n}-\boldsymbol{\mu}_{\text{f}}^{n-1}\right),$
and for the smoother covariance,
[…]
$=\mathbf{R}_{\text{s}}^{j,n-1}+\mathbf{D}^{j,n-2}\left(\mathbf{R}_{\text{s}}^{n-1,n}-\mathbf{R}_{\text{f}}^{n-1}\right),$
where by (B.3) we have (after the trivial reindexing $n-1\leadsto n-2$, without loss of generality),
$\mathbf{D}^{j,n-2}=\overset{{\curvearrowright}}{\prod^{n-2}_{i=j}}\mathbf{E}^{i}.$
With these we have proved (3.11) and (3.12) for $n\in\mathbb{N}$ and $j=0,1,\ldots,n-1$. For $j=n$, by the fundamental filter-smoother relation, we have $\boldsymbol{\mu}_{\text{s}}^{n,n}=\boldsymbol{\mu}_{\text{f}}^{n}$, $\mathbf{R}_{\text{s}}^{n,n}=\mathbf{R}_{\text{f}}^{n}$. This finishes the proof of Theorem 3.1. ∎
Appendix C Details of the Online Expectation-Maximization Algorithm for Parameter Estimation
In this appendix we outline the basic learning EM algorithm which sets the baseline structure for the online model identification algorithm developed in Section 4.3.1. As usual, it is assumed that the coupled CGNS in (2.1)–(2.2) is only partially observed, and based on the time discretization setup adopted thus far in this work (see (3.1)–(3.2) in Section 3.1), we denote the set of values of the state variables, when evaluated on the discrete-time steps $t_{j}$, as $\mathcal{X}=\{\mathbf{x}^{0},\ldots,\mathbf{x}^{j},\ldots,\mathbf{x}^{n}\}$ and $\mathcal{Y}=\{\mathbf{y}^{0},\ldots,\mathbf{y}^{j},\ldots,\mathbf{y}^{n}\}$, for the observable and unobservable components, respectively.
Given an ansatz of the model, the goal here is to maximize an objective function, specifically the log-likelihood of the observational process, with respect to the model parameters,
$\ell(\boldsymbol{\theta}):=\log\left(p\big{(}\mathcal{X};\boldsymbol{\theta}\big{)}\right)=\log\left(\int_{\mathcal{Y}}p\big{(}\mathcal{X},\mathcal{Y};\boldsymbol{\theta}\big{)}{\,\rm d}\mathcal{Y}\right),$
where $\boldsymbol{\theta}$ is the collection of model parameters. This is also known as the marginal log-likelihood of the observed data, with the full log-likelihood being given by $\log\left(p\big{(}\mathcal{X},\mathcal{Y};\boldsymbol{\theta}\big{)}\right)$. Using now any distribution over the discretized hidden variables, with density $q(\mathcal{Y})$, a lower bound on the marginal log-likelihood can be obtained in a rather trivial manner in the following way [136]:
$\ell(\boldsymbol{\theta})=\log\left(\int_{\mathcal{Y}}p\big{(}\mathcal{X},\mathcal{Y};\boldsymbol{\theta}\big{)}{\,\rm d}\mathcal{Y}\right)$ $=\log\left(\int_{\mathcal{Y}}q(\mathcal{Y})\frac{p\big{(}\mathcal{X},\mathcal{Y};\boldsymbol{\theta}\big{)}}{q(\mathcal{Y})}{\,\rm d}\mathcal{Y}\right)$
[…]
$=\int_{\mathcal{Y}}q(\mathcal{Y})\log\left(p\big{(}\mathcal{X},\mathcal{Y};\boldsymbol{\theta}\big{)}\right){\,\rm d}\mathcal{Y}-\int_{\mathcal{Y}}q(\mathcal{Y})\log(q(\mathcal{Y})){\,\rm d}\mathcal{Y}$
$:=\mathcal{F}(q,\boldsymbol{\theta}),$
where the negative value of $\int_{\mathcal{Y}}q(\mathcal{Y})\log\left(p\big{(}\mathcal{X},\mathcal{Y};\boldsymbol{\theta}\big{)}\right){\,\rm d}\mathcal{Y}$ is the so-called free energy, while $\mathcal{S}(q)=-\int_{\mathcal{Y}}q(\mathcal{Y})\log(q(\mathcal{Y})){\,\rm d}\mathcal{Y}$ is the Shannon differential entropy. Therefore, based on the fact that $\mathcal{F}(q,\boldsymbol{\theta})\leq\ell(\boldsymbol{\theta})$ for all distributions $q$ over $\mathcal{Y}$, it is clear that maximizing the marginal log-likelihood is equivalent to maximizing $\mathcal{F}$ alternatively with respect to the distribution $q$ and the parameters $\boldsymbol{\theta}$ [142, 143], thus creating a bridge between the usual MLE algorithm and EM procedure. This alternate maximization of $\mathcal{F}$ is the essence of the EM algorithm, where the EM designation stems from the following procedure:
  • E-Step:
[…]
  • M-Step:
    $\boldsymbol{\theta}_{n+1}\leftarrow\underset{\boldsymbol{\theta}}{\mathrm{argmax}}\left\{\mathcal{F}(q_{n+1},\boldsymbol{\theta})\right\}$
An essential observation in this process, is that the maximization in the E-Step is exactly reached when $q$ is the posterior conditional distribution of $\mathcal{Y}$, that is
$q_{n+1}(\mathcal{Y})=p(\mathcal{Y}|\mathcal{X};\boldsymbol{\theta}_{n}),$
which in the online setting of sequential arrival of observations corresponds to the online smoother posterior distribution. This is attributed to the equality condition for the Jensen inequality, which we used to prove the bound between the marginal log-likelihood and the objective function $\mathcal{F}$. In Jensen’s integral inequality, $\phi(\mathbb{E}\left[X\right])\leq\mathbb{E}\left[\phi(X)\right]$ for $\phi$ being operator convex and $X$ an arbitrary-valued random variable, it is known that equality holds if and only if $\phi$ is an affine function or if $X$ is constant almost surely. As such, since the objective function $\mathcal{F}$ can be rewritten in the following much more intuitive form:
$\mathcal{F}(q,\boldsymbol{\theta})=-\mathcal{P}(q,p(\boldsymbol{\cdot}|\mathcal{X};\boldsymbol{\theta}))+\ell(\boldsymbol{\theta}),$
where $\mathcal{P}$ is the relative entropy from Section 3.4.1, then equality holds when $\log\left(\frac{p\big{(}\mathcal{X},\mathcal{Y};\boldsymbol{\theta}\big{)}}{q(\mathcal{Y})}\right)$ is an affine function, which is exactly achieved when $q(\mathcal{Y})=p(\mathcal{Y}|\mathcal{X};\boldsymbol{\theta})\propto p\big{(}\mathcal{X},\mathcal{Y};\boldsymbol{\theta}\big{)}$, with the proportionality being due to Bayes’. Of course, this fact can also be seen through the Gibb’s inequality (see Section 3.4.1), since for $q(\mathcal{Y})=p(\mathcal{Y}|\mathcal{X};\boldsymbol{\theta})$ the relative entropy term on the right-hand side vanishes [144]. In such a situation, the bound $\mathcal{F}(q,\boldsymbol{\theta})\leq\ell(\boldsymbol{\theta})$ becomes an equality, $\mathcal{F}(q,\boldsymbol{\theta})=\ell(\boldsymbol{\theta})$. Note that the conditional distribution in the E-Step is very difficult to solve for general CTNDSs. Various numerical methods and approximations are often used [136, 135], which usually suffer from both approximation errors and the curse of dimensionality. Nevertheless, for the CGNS framework, the distribution $p(\mathcal{Y}|\mathcal{X};\boldsymbol{\theta}_{k})$ is given in an optimal and unbiased manner by the closed analytic formulae of the online discrete smoother from Theorem 3.1, which greatly facilitates the application of the online EM parameter estimation algorithm to many nonlinear models. On the other hand, since the differential entropy does not depend on $\boldsymbol{\theta}$, the maximum in the M-Step is obtained by maximizing the negative of the free energy,
$\boldsymbol{\theta}_{n+1}\leftarrow\underset{\boldsymbol{\theta}}{\mathrm{argmax}}\left\{\int_{\mathcal{Y}}p\big{(}\mathcal{Y}\big{|}\mathcal{X};\boldsymbol{\theta}_{n}\big{)}\log\left(p\big{(}\mathcal{X},\mathcal{Y};\boldsymbol{\theta}\big{)}\right){\,\rm d}\mathcal{Y}\right\}.$
This interpretation of the basic learning EM algorithm is known as the “maximization-maximization procedure” formulation, where the EM algorithm is viewed as two alternating maximization steps, which is a specific application of coordinate descent. We finally note that the EM algorithm enjoys a monotonicity property, which states that improving the negative of the free energy will at least not make the marginal log-likelihood worse with regards to the optimization goal [145], and that a convergence analysis for the EM method (including results for distributions outside the exponential family), can be found in the fundamental theoretical work of Wu [146]. As for the explicit expression of $\boldsymbol{\theta}_{n+1}$ in the case of the CGNS, as well as of the noise feedbacks in the case of unknown uncertainty matrices, both with respect to the posterior statistics, see the work of Chen [138].
Finally, here we provide a simple numerical trick for accelerating the convergence of the parameters to their true values in the online EM algorithm. This process is based on momentum-based methods from classical convex optimization, which utilize momentum terms into the update to aid the learning process (e.g., Nesterov’s accelerated gradient descent [147]). We denote by $\boldsymbol{\theta}_{n}$ and $\boldsymbol{\theta}_{n+1}$ the learned parameters (or part of them) in the previous ($n$) and current ($n+1$) iteration, respectively. An acceleration of the parameter value at the current step can be achieved by
$\boldsymbol{\theta}_{n+1}^{\text{new}}=\boldsymbol{\theta}_{n}+\alpha\left(\boldsymbol{\theta}_{n+1}-\boldsymbol{\theta}_{n}\right),$
where $\alpha\in(0,1]$ is a hyperparameter. If $\alpha=1$ then there is no acceleration; the acceleration rate depends on the amplitude of $\alpha$. Such a trick can be applied in the first few iterations of the online EM algorithm, especially for those in the unobserved process when the observability of the system is weak.